### v2

# General Distribution Learning: A theoretical framework for Deep Learning

Abstract
There remain numerous unanswered research questions on deep learning (DL) within the classical learning theory framework. These include the remarkable generalization capabilities of overparametrized neural networks (NNs), the efficient optimization performance despite non-convexity of objectives, the mechanism of flat minima for generalization, and the exceptional performance of deep architectures in solving physical problems. This paper introduces General Distribution Learning (GD Learning), a novel theoretical learning framework designed to address a comprehensive range of machine learning and statistical tasks, including classification, regression and parameter estimation. Departing from traditional statistical machine learning, GD Learning focuses on the true underlying distribution. In GD Learning, learning error, corresponding to the expected error in classical statistical learning framework, is divided into fitting errors due to models and algorithms, as well as sampling errors introduced by limited sampling data. The framework significantly incorporates prior knowledge, especially in scenarios characterized by data scarcity, thereby enhancing performance. Within the GD Learning framework, we demonstrate that the global optimal solutions in non-convex optimization can be approached by minimizing the gradient norm and the non-uniformity of the eigenvalues of the model’s Jacobian matrix. This insight leads to the development of the gradient structure control algorithm. GD Learning also offers fresh insights into the questions on deep learning, including overparameterization and non-convex optimization, bias-variance trade-off, and the mechanism of flat minima. The article structure is as follows: An initial overview of GD Learning is presented, outlining its fundamental components and a schematic diagram. We then delve into the components of the proposed framework, analyzing the standard loss function, gradient structure control algorithm, and establish an upper bound on learning error. Finally, we illustrate how GD Learning unifies classification and regression problems and addresses the unanswered questions on DL within the classical learning theory framework.
Keywords: distribution learning, deep learning, non-convex optimization
1 Introduction
Along with the remarkable practical achievements of deep learning, several fundamental questions on DL remain unanswered within the classical learning theory framework. These include: why deep neural networks (DNNs) can avoid overfitting despite their immense capacity, and how they surpass shallow architectures? how DNNs overcome the curse of dimensionality, and why optimization algorithms often succeed in finding satisfactory solutions despite non-convex non-linear, and often non-smooth DL objective? and which aspects of an architecture affect the performance of the associated models and the mechanisms behind these influences? Delving into the reasons and factors that govern the performance of DNNs not only enhances the interpretability for DNNs but also facilitates the development of more principled and reliable architecture designs.
Extensive empirical evidence indicates that, in addition to expressivity (Leshno et al., 1993; Barron, 1993) and trainability (Choromańska et al., 2014; Kawaguchi, 2016), the generalization of DNNs significantly influences prediction performance (Kawaguchi et al., 2017). Generalization error, a measure of model’s predictive ability on unseen data, is a critical aspect of learning models. Some classical theory work attributes generalization ability to the utilization of a low-complexity class of hypotheses and proposes various complexity measures to control generalization error. These include the Vapnik-Chervonenkis (VC) dimension (Wu, 2021), Rademacher complexity (Bartlett and Mendelson, 2003), and uniform stability (Mukherjee et al., 2006; Bousquet and Elisseeff, 2002). However, recent studies by (Zhang et al., 2016, 2021; He et al., 2015a; Belkin et al., 2018a) demonstrate that high-capacity DNNs can still generalize well with genuine labels, yet generalize poorly when trained with random labels. This apparent contradiction with the traditional understanding of overfitting in complex models highlights the need for further investigation (Vapnik and Alexey, 1971; Wu, 2021; Bartlett and Mendelson, 2003).
Distribution learning (estimation), a cornerstone in various fields such as machine learning, theoretical and applied statistics, information theory and communication applications, has a rich history. The distribution, described by a probability density/mass function (PDF/PMF), completely characterizes the “behavior” of a random variable, making distribution learning fundamental for investigating the properties of a given data set and for efficient data mining. The goal of distribution learning is to estimate the true underlying distribution from a given sample set. The majority of machine learning tasks can be framed with the distribution learning context, with both classification and regression problems mathematically represented by learning the conditional probability distributions between features and labels. In particular, generative learning corresponds to learning the underlying joint distribution. Consequently, approaching machine learning and deep learning from the perspective of distribution learning is a logical and coherent choice, given its inherent suitability for these tasks.
In this paper, we introduce a novel framework, General Distribution Learning (GD Learning). Departing from traditional statistical learning framework, GD Learning centers on the true underlying distribution, and employs norms to measure learning error. Learning error in this framework is divided into fitting errors caused by models and algorithms, as well as sampling errors due to limited sampling data. The GD Learning framework consists of a standard loss function, a fitting model, the prior knowledge, and a gradient structure control algorithm. The standard loss function, a constrained function, encompasses common loss functions and can be expressed as a Young–Fenchel loss function or a Bregman divergence. The fitting model in GD Learning is used to estimate the parameters of a distribution. It is important to note that when the distribution to be learned has finite supports, the support probabilities are the parameters to be estimated. The gradient structure control algorithm is designed to minimize the gradient norm and the structure error which is bounded by the non-uniformity of the model’s Jacobian eigenvalues. A key departure from traditional optimization algorithms is that gradient structure control algorithm optimizes for stationary points rather, not just global optima. Within the GD Learning framework, we demonstrate that the global optimal solution to non-convex optimization problems, such as fitting error minimization, can be approached by gradient structure control algorithm. GD Learning also offers a fresh perspective on deep learning issues, including overparameterization, non-convex optimization, the bias-variance trade-off, and the mechanisms behind flat minima.
The article is organized as follows: In Section 2, we delves into related work, reviewing learning theory, modern deep learning analysis, and distribution learning methods. In Section 3 we define the fundamental concepts and objects utilized in this paper. In Section 4 we provide an overview of GD Learning, detailing its basic components and schematic diagram. In Section 5, we define the standard loss and demonstrate its equivalence to both Young-Fenchel loss and Bregman divergence. In Section 6, we propose the gradient structure control algorithm and establish that the global optimal solution for non-convex optimization problems, such as fitting error minimization, can be approached by minimizing the gradient norm and the structure error which is bounded by the non-uniformity of the eigenvalues of the model’s Jacobian matrix. In Section 7, we derive the theoretical bounds for learning error, providing insights into the performance of GD Learning. In Section 8, we illustrate how GD Learning unifies classification and regression problems, and address the questions that remain unanswered within the classical learning theory framework. In Section 9, we summarize our findings and discuss their broader implications for the field.
2 Related work
In this section, we provide an overview of learning theory, modern approaches for analyzing deep learning, and distribution learning methods.
2.1 Learning theory
We begin by presenting a succinct summary of the classical mathematical and statistical theory underlying machine learning tasks and algorithms which, in their most general form, can be formulated as follows.
Definition 1 (Learning task)
In a learning task, one is given a loss function $\ell:\mathcal{M}(\mathcal{X},\mathcal{Y})\times\mathcal{Z}\to\mathbb{R}$ and training data $s^{n}=\{z^{(i)}\}_{i=1}^{n}$ which is generated by independent and identically distributed (i.i.d.) sampling according to the unknown true distribution $Z\sim\bar{q}$, where $Z$ are random variables which take the values in $\mathcal{Z}$. The objective is to construct a learning algorithm $\mathcal{A}$ that uses training data $s$ to find a model $f_{s}$ that performs well on the whole data, where $f_{s}=\mathcal{A}(s)\in\mathcal{F}$ , the hypothesis set $\mathcal{F}\subset\mathcal{M}(\mathcal{X},\mathcal{Y})$ , and the mapping $\mathcal{A}:s^{n}\to\mathcal{F}$. Here, performance is measured via $\mathbb{E}_{Z}\ell(f_{s},z)$.
The above definition can describe supervised learning task. It is noteworthy that unsupervised and semi-supervised learning tasks are also treated as supervised learning tasks when the training data comprises solely features $x$. Specifically, one can construct a label $y=T(x)$ where $x\in\mathcal{X},T(x)\in\mathcal{Y}$, with $T:\mathcal{X}\to\mathcal{Y}$ representing a transformation. Consequently, the learning task is to find a model $f_{s}$ approximating the transformation $T$, which is, often pursued to learn feature representations or invariances.
Given a hypothesis set, a widely utilized learning algorithm is empirical risk minimization (ERM), which minimizes the average loss on the given training data, as described in the subsequent definitions.
Definition 2 (Empirical risk)
For a given training dataset $s$ and a function $f\in\mathcal{M}(\mathcal{X},\mathcal{Y})$, the empirical risk is defined as:
^$\widehat{\mathcal{R}}_{s}(f):=\frac{1}{n}\sum_{i=1}^{n}\ell(f,z^{(i)}).$ (1)
Definition 3
For a function $f\in\mathcal{M}(\mathcal{X},\mathcal{Y})$, the expected risk is defined by:
$\mathcal{R}(f):=\mathbb{E}_{Z\sim\bar{q}}[\ell(f,z^{(i)})].$ (2)
Let $f_{s}=\mathcal{A}(s)$, then the expected risk of the model $f_{s}$ is given by
$\mathcal{R}(f_{s})=\mathbb{E}_{Z\sim\bar{q}}[\ell(f_{s},z^{(i)})].$ (3)
Definition 4 (ERM learning algorithm)
Given a hypothesis set $\mathcal{F}$, an Empirical Risk Minimization (ERM) algorithm $\mathcal{A}^{erm}$ chooses a function $\widehat{f}_{s}\in\mathcal{F}$ that minimizes the empirical risk for the training data $s$, i.e.,
$\mathcal{A}^{erm}(s)\in\arg\min_{f\in\mathcal{F}}\widehat{\mathcal{R}}_{s}(f).$ (4)
Definition 5 (Bayes-optimal function)
A function $f^{\star}\in\mathcal{M}(\mathcal{X},\mathcal{Y})$ achieving the minimum risk, its Bayes risk is
$\mathcal{R}^{\star}:=\inf_{f\in\mathcal{M}(\mathcal{X},\mathcal{Y})}\mathcal{R}(f),$ (5)
Definition 6 (Errors of learning)
Let $f^{\star}_{\mathcal{F}}\in\arg\min_{f\in\mathcal{F}}(f)$ represent the best approximation in $\mathcal{F}$, the optimization error denoted by $e_{\rm opt}$, generalization error denoted by $e_{\rm gen}$ and approximation error denoted by $e_{\rm approx}$ are defined as follows
$e_{\rm opt}(\mathcal{A},\mathcal{F}):=\widehat{\mathcal{R}}_{s}(f_{s})-\widehat{\mathcal{R}}_{s}(\widehat{f}_{s})\geq\widehat{\mathcal{R}}_{s}(f_{s})-\widehat{\mathcal{R}}_{s}(f^{\star}_{\mathcal{F}})$ (6)
$e_{\rm gen}(\mathcal{F}):=\sup_{f\in\mathcal{F}}|\mathcal{R}(f)-\widehat{\mathcal{R}}_{s}(f)|\geq\max\{|\mathcal{R}(f_{s})-\widehat{\mathcal{R}}_{s}(f_{s})|,|\mathcal{R}(f_{\mathcal{F}}^{\star})-\widehat{\mathcal{R}}_{s}(f^{\star}_{\mathcal{F}})|\}$
$e_{\rm approx}(\mathcal{F}):=\mathcal{R}(f^{\star}_{\mathcal{F}})-\mathcal{R}^{\star}.$
The error $\mathcal{R}(f_{s})-\mathcal{R}^{\star}$ can then be bounded by:
$\mathcal{R}(f_{s})-\mathcal{R}^{\star}$ $=\mathcal{R}(f_{s})-\widehat{\mathcal{R}}_{s}(f_{s})+\widehat{\mathcal{R}}_{s}(f_{s})-\widehat{\mathcal{R}}_{s}(f^{\star}_{\mathcal{F}})$ (7)
+^$+\widehat{\mathcal{R}}_{s}(f^{\star}_{\mathcal{F}})-\mathcal{R}_{s}(f^{\star}_{\mathcal{F}})+\mathcal{R}_{s}(f^{\star}_{\mathcal{F}})-\mathcal{R}^{\star}$
$\leq e_{\rm opt}(\mathcal{A},\mathcal{F})+2e_{\rm gen}(\mathcal{F})+e_{\rm approx}(\mathcal{F}).$
The optimization error $e_{\rm opt}$ is primarily influenced by the algorithm $\mathcal{A}$ that is used to find the model $f_{s}$ in the hypothesis set $\mathcal{F}$ for given training data $s$. By far the gradient-based methods are the most prevalent approach to address this issue, leveraging their ability to accurately and efficiently compute the derivatives by automatic differentiation in a lot of widely used models. Owing to the hierarchical structure of neural networks, the back-propagation algorithm has been proposed to efficiently calculate the gradient with respect to the loss function (Kelley, 1960; Dreyfus, 1962; Rumelhart et al., 1986; Griewank and Walther, 2000). Stochastic gradient descent (SGD) introduced by Robbins (1951) is a typical gradient-based method used to identify the unique minimum of a convex function (Kiefer and Wolfowitz, 1952). In scenarios where the objective function is non-convex, SGD generally converges to a stationary point, which may be either a local minimum, a local maximum, or a saddle point. Nonetheless, there are indeed conditions under which convergence to local minima is guaranteed (Nemirovski et al., 2008; Ge et al., 2015; Lee et al., 2016; Jentzen et al., 2018).
The universal approximation theorem (Lewicki and Marino, 2003; ichi Funahashi, 1989; Hornik et al., 1989; Leshno et al., 1993) states that neural networks with only 2 layers can approximate any continuous function defined on a compact set up to arbitrary precision. This implies that the approximation error, $e_{\rm approx}$, decreases as the model’s complexity increases. The generalization error, $e_{\rm gen}(\mathcal{F})$, is defined as the difference between the true risk and the empirical risk. To bounding the generalization error, Hoeffding’s inequality (Hoeffding, 1994) alongside the union bound directly provides the PAC (Probably Approximately Correct) bound (Valiant, 1984). The PAC bound is described as follows:
$\Pr[e_{\rm gen}(\mathcal{F})\leq\varepsilon(\frac{1}{\delta},\frac{1}{n})]\geq 1-\delta$ (8)
In simper terms, when the sample $s^{n}$ is chosen randomly, good generalization is ensured with a high probability, at least $1-\delta$, by minimizing $\varepsilon(\frac{1}{\delta},\frac{1}{n})$. The function $\varepsilon$ depends mainly on the number of samples $n$ and the confidence parameter $\delta$. The uniform convergence bound is a classical type of generalization bounds based on measures of complexity of the set $\mathcal{F}$, such as the VC-dimension (Wu, 2021) and the Rademacher complexity (Bartlett and Mendelson, 2003)). These bounds are applicable to all the hypotheses within $\mathcal{F}$ and are expressed mathematically as follows:
$\Pr[e_{\rm gen}(\mathcal{F})\leq\varepsilon(\frac{1}{\delta},\frac{1}{n},\mathcal{F})]\geq 1-\delta$ (9)
These bound,s characterized by the supremum operator $\sup_{f\in\mathcal{F}}$, can be seen as a worst-case analysis.
A potential drawback of the uniform convergence bounds is that they are independent of the learning algorithm, i.e., they do not take into account the way the hypothesis space is explored. To tackle this issue, algorithmic-dependent bounds have been proposed to take advantage of some particularities of the learning algorithm, such as its uniform stability (Bousquet and Elisseeff, 2002) or robustness (Xu and Mannor, 2010). In such instances, the bounds obtained hold for the hypothesis set $\mathcal{A}(s)\in\mathcal{F}$ which is learned with the algorithm $\mathcal{A}$ using the sample $s$. The form of such bounds is represented as:
$\Pr[e_{\rm gen}(\mathcal{A},\mathcal{F})\leq\varepsilon(\frac{1}{\delta},\frac{1}{n},\mathcal{A})]\geq 1-\delta$ (10)
For instance, the work presented by Hardt et al. (2015) utilizes this approach to derive generalization bounds for hypotheses learned by stochastic gradient descent.
The classical trade-off highlights the need to balance model complexity and generalization. While simpler models may suffer from high bias, overly complex ones can overfit noises (high variance). However, reconciling the classical bias-variance trade-off with the empirical observation of good generalization, despite achieving zero empirical risk on noisy data using DNNs with a regression loss, presents a significant challenge (Belkin et al., 2018b). Generally speaking, in Deep learning, overparameterization and perfectly fitting noisy data does not exclude good generalization performance. Moreover, empirical evidence across various modeling techniques, including DNNs, decision trees, random features, linear models, indicates that test error can decreases even below the sweet-spot in the u-shaped bias-variance curve when further increasing the number of parameters (Belkin et al., 2018a; Geiger et al., 2019; Nakkiran et al., 2019). This phenomenon is often referred to as the double descent curve or benign overfitting, as illustrated in Figure 1.
Figure 1: The double descent risk curve, which incorporates the U-shaped risk curve (i.e., the “classical” regime) together with the observed behavior from using high capacity function classes (i.e., the “modern” interpolating regime), separated by the interpolation threshold. The predictors to the right of the interpolation threshold have zero training risk (Belkin et al., 2018a).
The available body of evidence suggests that classical tools from statistical learning theory alone, such as uniform convergence, algorithmic stability, or bias-variance trade-off may be unable to fully capture the behavior of DNNs (Zhang et al., 2016; Nagarajan and Kolter, 2019).
2.2 Deep learning theory
Deep learning poses several mysteries within the conventional learning theory framework, including the outstanding generalization ability of overparametrized neural networks, the impact of depth in deep architectures, the apparent absence of the curse of dimensionality, the surprisingly successful optimization performance despite the non-convexity of the problem, understanding what features are learned, why deep architectures perform exceptionally well in solving physical problems, and which fine aspects of an architecture affect the behavior of a learning task in which way. Theoretical insights into these problems from prior research have predominantly followed two main avenues: architecture of model and optimization algorithm.
Regarding model architecture, Hanin and Sellke (2017); Kidger and Lyons (2019) show that very narrow neural networks can remain universally approximative with an appropriate selection of depth. More specifically, an increase in the depth of a neural network architecture has the effect of increasing its expressiveness exponentially (Poole et al., 2016; Zhang et al., 2023). Increasing the depth of a neural network architecture naturally yields a highly overparametrized model, whose loss landscape is known for a proliferation of local minima and saddle points (Dauphin et al., 2014). Overparameterization is widely believed to be crucial for the generalization ability of DNNs, as supported by studies like (Du et al., 2018; Yun et al., 2018; Chizat et al., 2018; Arjevani and Field, 2022). It is also widely believed that the generalization of highly overparametrized models cannot be adequately explained without accounting for stochastic optimization algorithms whose inherent stochasticity performs implicit regularization, guiding parameters towards flat plateaus with lower generalization errors (Keskar et al., 2016; Chaudhari and Soatto, 2017). First-order optimization methods, such as SGD, have been demonstrated to be more efficient in escaping from saddle points and converge towards to flat minima (Dauphin et al., 2014; Lee et al., 2019). To identify the flat minima, numerous mathematical definitions of flatness exist, with a prevalent one being the largest eigenvalue of the Hessian matrix of the loss function with respect to the model parameters on the training set, denoted as $\lambda_{\max}$ (Jastrzebski et al., 2017; Lewkowycz et al., 2020; Dinh et al., 2017). The work of Jacot et al. (2018) elucidates that the evolution of the trainable parameters in continuous-width DNNs during training can be captured by the neural tangent kernel (NTK). Recent studies demonstrate that with specialized scaling and random initialization, the dynamics of gradient descent (GD) in continuous-width multi-layer neural networks can be tracked via NTK and the convergence to the global optimum with a linear rate can be proved (Du et al., 2018; Zou et al., 2018; Allen-Zhu et al., 2018) . An alternative research direction attempts to examine the infinite-width neural network from a mean-field perspective (Sirignano and Spiliopoulos, 2018; Mei et al., 2018; Chizat and Bach, 2018). The key idea is to characterize the learning dynamics of noisy SGD as the gradient flows over the space of the probability distribution of the neural network parameters. NTK and mean-filed methods provide a fundamental understanding on training dynamics and convergence property, despite the analysis is seldom connected to the deep representation viewpoint. Nonetheless, from the perspective of objective function optimization, training DNNs is a non-convex optimization problem, yet empirical observations indicates that even simple algorithms, such as SGD, tend to converge to the global minimum. The mechanism behind this cannot be explained by classical convex optimization theory. A popular claim in the literature is that the overparameterization also plays an important role in non-convex optimization (Du et al., 2018; Zou et al., 2018; Allen-Zhu et al., 2018; Sirignano and Spiliopoulos, 2018; Mei et al., 2018; Chizat and Bach, 2018).
Accumulating evidence indicates that the exceptional performance of DNNs is attributed to a confluence of factors, including their architectural design, optimization algorithms, and even training data.
2.3 Distribution learning
In statistical data analysis, determining the probability density function (PDF) or probability mass function (PMF) from a limited number of random samples is a common task, especially when dealing with complex systems where accurate modeling is challenging. Generally, there are three main approaches to this problem: parametric, semi-parametric and non-parametric models.
Parametric models assume the data follows a specific distribution with unknown parameters. Examples include Gaussian model, Gaussian Mixture models (GMMs), and distributions from the generalized exponential family. A major limitation of these methods is that they apply quite simple parametric assumptions, which may not be sufficient or verifiable in practise. For instance, GMMs are based on the first few moments, and thus provide a poor fit for some distributions, including heavy-tailed processes (Rahman et al., 2008).
Non-parametric models estimate the PDF/PMF based on data points without making assumptions about the underlying distribution (Härdle et al., 2004). Examples include histogram (Sturges, 1926; Scott, 1979; Freedman and Diaconis, 1981) and kernel density estimation (KDE) (Parzen, 1962; Silverman, 2018; Hall et al., 2004). While these methods are more flexible, they may suffer from overfitting or poor performance in high-dimensional spaces due to the curse of dimensionality.
Semi-parametric models (McLachlan et al., 2019) strike a balance between parametric and non-parametric approaches. They do not assume the a prior-shape of the PDF/PMF to estimate as non-parametric techniques. However, unlike the non-parametric methods, the complexity of the model is fixed in advance, in order to avoid a prohibitive increase of the number of parameters with the size of the data set, and to limit the risk of overfitting. Finite mixture models are commonly used to serve this purpose.
Statistical methods for determining the parameters of model conditional on the data, including standard likelihood method (Aldrich, 1997; Hald, 1999; Meeker et al., 2022), M-estimators (Huber, 1964, 1973), Bayesian methods (MacKay, 2003; Sivia and Skilling, 2006; Murphy, 2022), have been widely employed. The Maximum Likelihood (MLE) introduced by Fisher (Aldrich, 1997), is commonly used in data regression, with earlier ideas discussed by Daniel Bernoulli and and Gauss (Hald, 1999). In 1964, Huber (1964) proposed a generalization of the maximum likelihood objective function. In 1973 Huber (1973) extended the idea of using M-estimators for solution of regression problems. However, traditional estimators struggle in high-dimensional space: because the required sample size or computational complexity often explodes geometrically with increasing dimensionality, it has almost no practical value for high-dimensional data.
Recently, neural network-based approaches are proposed for PDF/PMF estimation and yield promising results in problems with high-dimensional data points such as images. There are mainly two families of such neural density estimators: auto-regressive models (Larochelle and Murray, 2011; Papamakarios et al., 2017) and normalizing flows (Rezende and Mohamed, 2015; Dinh et al., 2016). Autoregression-based neural density estimators decompose the density into the product of conditional densities based on probability chain rule $p(x)=\prod_{i}p(x_{i}|x_{1:i-1})$. Each conditional probability $p(x_{i}|x_{1:i-1})$ is modeled by a parametric density (e.g., Gaussian or mixture of Gaussian), of which the parameters are learned by neural networks. Density estimators based on normalizing flows represent $x$ as an invertible transformation of a latent variable $z$ with known density, where the invertible transformation is a composition of a series of simple functions whose Jacobian is easy to compute. The parameters of these component functions are then learned by neural networks As suggested in Kingma et al. (2016), both of these are special cases of the following general framework. Given a differentiable and invertible mapping $G:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and a base density $p_{\mathbf{z}}(\mathbf{z})$, the density of $\mathbf{x}=G(\mathbf{z})$ can be represented using the change of variable rule as follows:
$p_{\mathbf{x}}(\mathbf{x})=p_{\mathbf{z}}(\mathbf{z})|\mathrm{det}(\mathbf{J}_{\mathbf{z}})|^{-1}$ (11)
where $\mathbf{J}_{\mathbf{z}}=(\partial G(\mathbf{z}))/\partial\mathbf{z}^{\top}$ is the Jacobian matrix of function $G(\cdot)$ at point $\mathbf{z}$. Density estimation at $\mathbf{x}$ can be solved if the base density $p_{\mathbf{z}}(\mathbf{z})$ is known and the determinant of Jacobian matrix is feasible to calculate. To achieve this, previous neural density estimators have to impose heavy constraints on the model architecture. For example Papamakarios et al. (2017); Dinh et al. (2016); Kingma et al. (2016) require the Jacobian to be triangular.
3 Preliminaries
3.1 Notation
  • The set of real numbers is denoted by $\mathbb{R}$, and the cardinality of set $\mathcal{Z}$ is denoted by $|\mathcal{Z}|$.
  • The $|\mathcal{Z}|$-dimensional probability simplex is denoted by $\Delta^{|\mathcal{Z}|}:=\{p\in\mathbb{R}_{+}^{|\mathcal{Z}|}\colon\|p\|_{1}=1\}$.
  • Random variables are denoted using upper case letters such as $X,Y$ which take the values in set $\mathcal{X}$, and $\mathcal{Y}$, respectively.
  • The expectation of a random variable $X$ is denoted by $\mathbb{E}{[X]}$ or $\mathbb{E}_{X}[X]$, with $\mathbb{E}_{X\sim p}[X]$ when the distribution of $X$ is specified.
  • The space of fitting models, which is a set of functions endowed with some structure, is represented by $\mathcal{F}({\Theta})=\{f(\theta):\theta\in\Theta\}$, where $\Theta$ denotes the parameter space.
  • Balls with respect to various norms are denoted as $\mathcal{B}_{q}(r):=\{x\in\mathbb{R}^{m},\|x\|_{q}\leq r\}$.
  • The Fréchet sub-differential is defined as $\partial f(x)=\{s\in\mathbb{R}^{n}|\lim_{y\to x}\frac{f(y)-f(x)-\langle s,y-s\rangle}{\|y-x\|}\geq 0\}$ (see e.g., (Li et al., 2020),Page 27). If $f(x)$ is smooth and convex, the Fréchet sub-differential ^$\hat{\partial}f(x)$ reduces to the usual gradient, i.e., $\partial f(x)=\{\nabla f(x)\}$.
  • The Fenchel conjugate of a function $\Omega$ is denoted by $\Omega^{*}(f):=\sup_{p\in\mathrm{dom}(\Omega)}\langle p,f\rangle-\Omega(p)$. By default, $\Omega$ is a continuous strictly convex function, and its gradient with respect to $p$ is denoted by $p_{\Omega}^{*}$.
  • The Bregman divergence associated with a convex function $\Omega$ and its subgradient $f$ is defined as $B_{\Omega}(q,p)=\Omega(q)-\Omega(p)-f^{\top}(q-p)$, where $f\in\partial\Omega(p)$.
  • The Kullback-Leibler (KL) divergence is denoted by
    $D_{KL}(p\|f):=\sum_{i}p_{i}\log\frac{p_{i}}{f_{i}}-\sum_{i}p_{i}+\sum_{i}f_{i}.$ (12)
  • The largest and smallest eigenvalues of matrix $A$ are denoted by $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$, respectively.
3.2 Setting and basics
The Random pair $Z=(X,Y)$ follows the distribution $\bar{q}$ and take the values in the product space $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ representing the input feature space and $\mathcal{Y}$ the label set. The training data, denoted by an $n$-tuple $s^{n}=\{z{(i)}\}_{i=1}^{n}=\{(x^{(i)},y^{(i)})\}_{i=1}^{n}$, is composed of i.i.d. samples drawn from the unknown true distribution $\bar{q}$. The empirical estimator $q^{n}$ (or simply $q$) is derived from the observed dataset $s^{n}$ and prior knowledge, functioning as an unbiased estimate of $\bar{q}$. $q^{n}$ obeys the (weak) law of large numbers, ensuring convergence to $\bar{q}$ in probability. In the absence of prior knowledge about $\bar{q}$, the Monte Carlo method (MC), which estimates the distribution based on the event frequencies, denoted by $\hat{q}(z):=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{z\}}z^{(i)}$, is commonly employed. In this article, we refer to $\hat{q}$ as the frequency distribution to distinguish it from $q$. Given that $\bar{q}$ is the probability density distribution function, we denote the empirical probability density distribution as $\hat{q}(z):=\frac{1}{n}\sum_{i=1}^{n}\delta_{z}$, where $\delta_{z}$ represents the Dirac delta function. Real-world distributions are often smooth, motivating the common practice of smoothing $\hat{q}$ to approximate the true distribution $\bar{q}$. These smoothing techniques effectively integrate prior knowledge and often advantageous, as exemplified by classical Laplacian and kernel density estimation (KDE) methods. These methods have demonstrated superior performance in numerous instances.
We employ $q_{X}(x)$ and $q_{Y|X}(y|x)$ to represent the marginal and conditional PMFs, respectively. For notational simplicity, we define
$q_{Y|x}(y):=q_{Y|X}(y|x),$ (13)
$q_{Y|x}:=(q_{Y|x}(y^{(1)}),\cdots,q_{Y|x}(y^{(|\mathcal{Y}|)}))^{\top},$
where $y^{(i)}\in\mathcal{Y}$ for each $i$.
Based on the Fenchel-Young inequality, it is natural to obtain the Fenchel-Young loss which unifies many existing losses, including the hinge, logistic, and sparse-max losses, as demonstrated by Blondel et al. (2019).
Definition 7 (Fenchel-Young loss generated by $\Omega$)
The Fenchel-Young loss $d_{\Omega}\colon\mathrm{dom}(\Omega)\times\mathrm{dom}(\Omega^{*})\to\mathbb{R}_{+}$ generated by $\Omega$ is defined as:
$d_{\Omega}(p;f):=\Omega(p)+\Omega^{*}(f)-\langle p,f\rangle,$ (14)
where $\Omega^{*}$ denotes the conjugate of $\Omega$.
Lemma 8 (Properties of Fenchel-Young losses)
The following are the properties of Fenchel-Young losses.
  1. 1.
    $d_{\Omega}(p,f)\geq 0$ for any $p\in\mathrm{dom}(\Omega)$ and $f\in\mathrm{dom}(\Omega^{*})$. If $\Omega$ is a lower semi-continuous proper convex function, then the loss is zero iff $f\in\partial\Omega(p)$. Furthermore, when $\Omega$ is strictly convex, the loss is zero iff $f=p_{\Omega}^{*}$.
  2. 2.
    $d_{\Omega}(p,f)$ is convex with respect to $p$ and its subgradients include the residual vectors: $p_{\Omega}^{*}-f\in\partial_{p}d_{\Omega}(p;f)$. If $\Omega$ is strictly convex, then $\nabla_{p}d_{\Omega}(p;f)=p_{\Omega}^{*}-f$.
  3. 3.
    The relation between Fenchel-Young losses and Bregman divergences can be further clarified using duality. Letting $f=\nabla\Omega(p)$ (i.e., $(p,f)$ is a dual pair), it follows that $B_{\Omega}(q\|p)=d_{\Omega}(q,f)$. In other words, Fenchel-Young losses can be viewed as a “mixed-form Bregman divergence” (Amari, 2016, Theorem 1.1 ), where the argument $p$ in Bregman divergence is replaced by its dual $f$ in Fenchel-Young loss.
Throughout this paper, being consistent with practical applications and without affecting the conclusions, we assume that $\Omega$ is a proper, lower semi-continuous (l.s.c.), and strictly convex function or functional. Consequently, the Fenchel-Young loss attains zero if and only if $f=p_{\Omega}^{*}$.
The theorems in the subsequent sections rely on the following lemmas.
Lemma 9 (Sanov’s theorem(Cover, 1999))
$\lim_{n\to\infty}-\frac{1}{n}\log\Pr(q|\bar{q})=D_{KL}(q\|\bar{q})$ (15)
Lemma 10 (Theorem 2(Bai and Yin, 1993))
Given a double array of i.i.d. random variables $\left[X_{uv};u,v=1,2,\ldots\right]$ with zero mean and unit variance, let matrix $X$ and $S$ be defined as:
$X=\left[X_{uv}:u=1,\ldots,p;v=1,\ldots,n\right],\quad S=(1/n)XX^{\prime}.$
Then, if $E\left|X_{11}\right|^{4}<\infty$, as $n\rightarrow\infty,p\rightarrow\infty,p/n\rightarrow y\in(0,1)$, almost surely,
$\lim\lambda_{\min}(S)=(1-\sqrt{y})^{2},$ (16)
$\lim\lambda_{\max}(S)=(1+\sqrt{y})^{2}.$
Lemma 11 (Markov inequality)
Suppose a random variable $Z\geq 0$, then for all $t\geq 0$, $\operatorname{Pr}(Z\geq t)\leq\mathbb{E}[Z]/t$.
4 Overview of GD Learning
In this section, we provide an in-depth overview of GD Learning, and compare it with statistical learning framework to highlight its distinct characteristics.
Figure 2: Basic components of GD Learning.
Figure 3: Schematic diagram of GD Learning.
GD Learning aims to estimate the underlying distribution $\bar{q}$, either $\bar{q}_{Y|X}$ or $\bar{q}_{X}$ using an i.i.d. sample set $s$. Therefore we formulate GD Learning into two situations:
  • native distribution learning
    $\min_{f\in\mathcal{F}}\mathbb{E}_{Z\sim\bar{q}}\|\bar{q}-p\|\,\ \mathrm{s.t.}\,\ s=\{z_{i}\}_{i=1}^{n},$ (17)
    where $p=M(f)$ with $M:\mathcal{F}\to\Delta$, denotes a mapping informed by prior knowledge about distribution $\bar{q}$ to convert parameter $f$ into a probability distribution. When $\bar{q}$ is a PMF over finite support, we set $M(f)=(f)_{\Omega^{*}}^{*}$.
  • conditional distribution learning
    $\min_{f_{x}\in\mathcal{F}}\mathbb{E}_{X\sim\bar{q}_{X}}\|\bar{q}_{Y|x}-p_{Y|x}\|\,\ \mathrm{s.t.}\,\ s=\{z_{i}\}_{i=1}^{n},$ (18)
    where $p_{Y|x}=M(f_{x})$, and $M$ similarly transforms $f_{x}$ to the corresponding probability distribution. Analogously, when $\bar{q}_{Y|x}$ is a PMF with finite support, $M(f_{x})=(f_{x})_{\Omega^{*}}^{*}$ is adopted.
Our work primarily employs $L_{2}$-norm minimization as the default optimization objective. It is justified to use any norm in finite-dimensional spaces, as the equivalence of different norms ensures that the conclusions derived are invariant to the choice of norm.
Conditional distribution learning is inherently linked to supervised learning. Native distribution learning, on the other hand, aligns with unsupervised learning, as it can be considered a specific case of conditional distribution learning when the input variable $X$ is constant, i.e., $X\sim\delta_{x}$. Thus, our primary focus is conditional distribution learning, and the insights derived can be extended to native distribution learning as well.
We introduce the discrepancy function
$R_{q_{X}}(p_{Y|x},p^{\prime}_{Y|x})=\mathbb{E}_{X\sim q_{X}}\|p_{Y|x}-p^{\prime}_{Y|x}\|_{2},$ (19)
which measures the average discrepancy between the conditional distribution $p_{Y|x}$ and $p^{\prime}_{Y|x}$ when the random parameter $X$ is sampled from the distribution $q_{X}$. Then the errors in GD Learning are defined as follows:
  1. 1.
    Learning Error: This error serves as the optimization objective, measuring how well the model approximates the underlying true distribution, denoted by $\|\bar{q}-p\|_{2}$. In conditional distribution learning, we extend this concept to the expected learning error, which is typically denoted as
    $R_{\bar{q}_{X}}(p_{Y|x},q_{Y|x})=\mathbb{E}_{X\sim\bar{q}_{X}}\|\bar{q}_{Y|x}-p_{Y|x}\|_{2},$ (20)
    where $\theta$ are the model’s parameters. From the definition, it is evident that learning error corresponds to expected risk in statistical learning framework.
  2. 2.
    Fitting Error: This error measures the model’s performance in fitting the estimator $q$, denoted by $\|q-p\|_{2}$. Similarly, in conditional distribution learning, we define the equivalent concept, the average fitting error, as follows:
    $R_{\hat{q}_{X}}(q_{Y|x},p_{Y|x})=\mathbb{E}_{X\sim\hat{q}_{X}}\|q_{Y|x}-p_{Y|x}\|_{2}.$ (21)
  3. 3.
    Sampling Error: This error, denoted by $\|q-\bar{q}\|_{2}$, arises from the finite sample size and serves as a model independent metric that characterizes the sampled data. For conditional distribution learning, the expected sampling error is defined as follows:
    $R_{\bar{q}_{X}}(q_{Y|x},\bar{q}_{Y|x})=\mathbb{E}_{X\sim\bar{q}_{X}}\|\bar{q}_{Y|x}-q_{Y|x}\|_{2}.$ (22)
As shown in figure 2, GD Learning’s architecture involves four main components: fitting model, standard loss, prior knowledge, and a gradient structure control algorithm (represented by $\mathcal{A}$) . The fitting model is a parametric function. It is generally assumed that the fitting model is a universal approximator including NN, Gradient Boosting Decision Tree (GBDT), Random Forest (RF), Support Vector Machine (SVM), XGBoost, etc. NNs, especially DNNs, are widely used due to their strong fitting ability and good generalization performance. Regarding the fitting ability of NN,the Universal Approximation Theorem (Hornik et al. (1989); Cybenko (1989)) states that a feedforward network with any “squashing” activation function $\sigma(\cdot)$, such as the logistic sigmoid function, can approximate any Borel measurable function $f(x)$ with any desired non-zero amount of error, provided that the network is given enough hidden layer size $n$. Universal Approximation theorems have also been proved for a wider class of activation functions, which includes the now commonly used ReLU (Leshno et al. (1993)). Generally speaking, having more parameters typically suggests a model has stronger fitting ability. However, this description is an oversimplification. The estimator $q$ of $\bar{q}$ is generated by combining prior knowledge and sampling instances. To address the appropriate utilization of prior knowledge with a given sample set, we introduce the concept of uncertainty. A comprehensive examination of this concept is presented in Section7. As for the standard loss function and gradient structure control algorithms, which differ from those the statistical learning framework, we analyze them in a separate usage section later in this paper. For a more transparent elucidation of the algorithm principle, we employ native distribution learning as an illustrative example. As shown in Figure 3, in GD Learning, $q$ serves as a bridge connecting the unknown true distribution $\bar{q}$ and the model output $p$. The optimization objective $\|p-\bar{q}\|$ can be bounded using the triangular inequality, decomposing it into the fitting error and the sampling error. Consequently, minimizing the sampling and fitting errors can effectively controls the learning error. A central aspect of sampling error is the selection of an unbiased estimator $q$ for $\bar{q}$. Although $\hat{q}(z)$ is theoretically sound, in practical scenarios, especially in the case of insufficient samples, techniques like Laplacian smoothing and kernel density estimation (KDE) often yield superior performance. To address this, we propose the principle of minimum uncertainty and utilize it to select $q$. Fitting error optimization is often a non-convex problem, and gradient descent algorithms do not guarantee a global optimal solution. To address this, we propose gradient structure control algorithm, and demonstrates that by reducing the gradient norm and the distance between the maximum and minimum eigenvalues of the Jacobian matrix of parameters, the parameters can approximate the global optimum.
5 Standard loss function
In this section, we formally introduce the concept of standard loss function and demonstrate its equivalence to both the Fenchel Young loss and Bregman divergence. We then delve into a detailed examination of its fundamental properties. Based on the standard loss function, we introduce the concept of structure error and derive the standard error equality which serves as a foundation for non-convex optimization in DL learning.
Definition 12 (Standard loss)
A loss function $\ell(q,p):\Delta\times\Delta\to\mathbb{R}^{+}$ is considered as a standard loss if it satisfies the following conditions:
  • Non-negativity and strict convexity: The loss is non-negative and strictly positive when $q\neq p$, ensuring that it is uniquely minimized when predicting the correct probability $q$.
  • Minimization with respect to the true distribution: The expected loss, $\mathbb{E}_{q}\ell(q,p)$, achieves its minimum value when $q=\bar{q}$, expressed as $\bar{q}=\arg\min_{p}\mathbb{E}_{q}\ell(q,p)$.
The uniqueness of the optimal solution allows for a clear connection between the standard loss and the optimization, as it guides the model towards the correct distribution. We establish the equivalence of the standard losses to the Fenchel-Young loss and Bregman divergence in the following theorem.
Theorem 13 (Uniqueness of standard loss)
Given any standard loss $\ell(q,p)$, there exists a strictly convex function $\Omega$ such that $\ell(q,p)=d_{\Omega}(q,p_{\Omega}^{*})=B_{\Omega}(q,p)$, where $d_{\Omega}(q,p_{\Omega}^{*})$ denotes the Fenchel-Young loss, and $B_{\Omega}(q,p)$ represents the Bregman divergence.
The proof of this result is provided in AppendixA.1.
In our study, we represent the standard loss as $d_{\Omega}(q,f_{\theta})$, where $f_{\theta}$ denotes the model parameterized by $\theta$. To maintain generality without compromising clarity, we assume that $d_{\Omega}(q,f_{\theta})$ is continuous and Fréchet differentiable with respect to $\theta$. The standard loss function is a broad concept that encompasses a wide range of commonly used loss functions in statistics and machine learning, as illustrated in Table1.
Table 1: Examples of regularized prediction functions and their corresponding standard losses(Blondel et al., 2019)
Loss $\mathrm{dom}(\Omega)$ $\Omega(\mu)$ $\hat{y}_{\Omega}(\theta)$ $d_{\Omega}(\theta;y)$
Squared $\mathbb{R}^{d}$ $\frac{1}{2}\|\mu\|^{2}$ $\theta$ $\frac{1}{2}\|y-\theta\|^{2}$
Perceptron $\Delta^{|\mathcal{Y}|}$ 00 $\arg\max(\theta)$ $\max_{i}\theta_{i}-\theta_{k}$
Logistic $\Delta^{|\mathcal{Y}|}$ $-H(\mu)$ $\mathrm{softmax}(\theta)$ $\log\sum_{i}\exp\theta_{i}-\theta_{k}$
Hinge $\Delta^{|\mathcal{Y}|}$ $\langle{\mu},{\mathbf{e}_{k}-\mathbf{1}}\rangle$ $\arg\max(\mathbf{1}-\mathbf{e}_{k}+\theta)$ $\max_{i}\leavevmode\nobreak\ [[i\neq k]]+\theta_{i}-\theta_{k}$
Sparsemax $\Delta^{|\mathcal{Y}|}$ $\frac{1}{2}\|\mu\|^{2}$ $\mathrm{sparsemax}(\theta)$ $\frac{1}{2}\|y-\theta\|^{2}-\frac{1}{2}\|hat{y}_{\Omega}(\theta)-\theta\|^{2}$
Logistic (one-vs-all) $[0,1]^{|\mathcal{Y}|}$ $-\sum_{i}H([\mu_{i},1-\mu_{i}])$ $\mathrm{sigmoid}(\theta)$ $\sum_{i}\log(1+\exp(-(2y_{i}-1)\theta_{i}))$
Structured perceptron $\mathrm{conv}(\mathcal{Y})$ 00 $\mathrm{MAP}(\theta)$ $\max_{y^{\prime}}\langle{\theta},{y^{\prime}}\rangle-\langle{\theta},{y}\rangle$
Structured hinge $\mathrm{conv}(\mathcal{Y})$ $-\langle{\mu},{c_{y}}\rangle$ $\mathrm{MAP}(\theta+c_{y})$ $\max_{y^{\prime}}\langle{c_{y}},{y^{\prime}}\rangle+\langle{\theta},{y^{\prime}}\rangle-\langle{\theta},{y}\rangle$
CRF $\mathrm{conv}(\mathcal{Y})$   $\min\limits_{p\in\Delta\colon\mathbb{E}_{p}[Y]=\mu}-H(p)$ $\mathrm{marginals}(\theta)$ $\log\sum_{y^{\prime}}\exp\langle{\theta},{y^{\prime}}\rangle-\langle{\theta},{y}\rangle$
SparseMAP $\mathrm{conv}(\mathcal{Y})$ $\frac{1}{2}\|\mu\|^{2}$ $\mathrm{sparsemap}(\theta)$ $\frac{1}{2}\|y-\theta\|^{2}-\frac{1}{2}\|\hat{y}_{\Omega}(\theta)-\theta\|^{2}$
  • 1
    For multi-class classification, we assume $\mathcal{Y}=\{\mathbf{e}_{i}\}_{i=1}^{d}$ and the ground-truth is $y=\mathbf{e}_{k}$, where $\mathbf{e}_{i}$ denotes a standard basis (“one-hot”) vector.
  • 2
    For structured classification, we assume that elements of $\mathcal{Y}$ are $d$-dimensional binary vectors with $d\ll|\mathcal{Y}|$, and we denote by $\mathrm{conv}(\mathcal{Y})=\{\mathbb{E}_{p}[Y]\colon p\in\Delta^{\mathcal{Y}}\}$ the corresponding marginal polytope (Wainwright and Jordan, 2008).
  • 3
    We denote by $H(p):=-\sum_{i}p_{i}\log p_{i}$ the Shannon entropy of a distribution $p\in\Delta^{\mathcal{Y}}$ and $\hat{y}_{\Omega}\in\arg\min_{\mu\in\mathrm{\Omega}}d_{\Omega}(\theta,\mu)$.
Based on the standard loss function, we introduce a novel metric, the structure error, to quantify the directional discrepancy between vectors $e$ and $g$ using the following formulation:
${T(e,g)}=\frac{1}{2}\min_{a\in\mathbb{R}^{+}}\{\|ae-\frac{g}{a}\|_{2}^{2},\|ae+\frac{g}{a}\|_{2}^{2}\},$ (23)
where $T(e,g)=0$ iif $\forall i,j,e(i)g(j)=e(j)g(i)$ holds. Notably, $T(e,g)$ is an even function with respect to $g$ by definition, that is, $T(e,g)=T(e,-g)$. When $e$ and $g$ are parameterized by the random variable $X\sim q_{X}$, we define the expected structure error as:
$\mathbb{E}_{X}T(e_{x},g_{x})=\frac{1}{2}\mathbb{E}_{X}\min_{a\in\mathbb{R}^{+}}$$\{\|ae_{x}-f_{x}/a\|_{2}^{2}$ (24)
$,\|ae_{x}+f_{x}/a\|_{2}^{2}\},$
and $\mathbb{E}_{X}T(e_{x},g_{x})=0$ iif $\forall i,j,x,g_{x}(i)e_{x}(j)=g_{x}(j)e_{x}(i)$ holds.
We present a theorem that offers a more readily optimized representation of the structure error, as detailed below.
Theorem 14 (structure error equality)
Given functions $e_{x},g_{x}$ parameterized by a random variable $X\sim q_{X}$, the following holds:
$T(x,f)$ $=\|f\|_{2}\|x\|_{2}-|\langle x,f\rangle|\geq 0,$ (25)
$\mathbb{E}_{X}T(e_{x},g_{x})$ $=\mathbb{E}_{X}\|e_{x}\|_{2}\|g_{x}\|_{2}-\mathbb{E}_{X}|\langle e_{x},g_{x}\rangle|$
$\geq 0.$
The proof of this result is provided in Appendix A.2.
6 Gradient structure control algorithm
In this section, we propose a novel gradient structure control algorithm to tackle the challenges faced by gradient-based optimization methods with non-convex objectives. Specifically, we demonstrate within the GD Learning framework, reducing the gradient norm and the discrepancy between the maximum and minimum eigenvalues of the Jacobian matrix of parameters, ensures convergence to the global optimum. We commence by defining an optimization problem of interest:
$\min_{\theta}d_{\Omega}(h,f_{\theta}),$ (26)
[…]
Introducing $A:=\nabla_{\theta}f^{\top}\nabla_{\theta}f$ as the matrix constructed from the product of the gradient of the function $f$ with respect to the parameters $\theta$ with itself, and $e:=\nabla_{f}d_{\Omega}(h,f)$ representing the gradient of $d_{\Omega}(h,f)$ with respect to $f$, we proceed to present a pivotal theoretical assertion:
Theorem 15 (bound of gradient structure control)
Assuming $\|Ae\|\neq 0$, the following inequality holds:
$\frac{\|\nabla_{\theta}d_{\Omega}(h,f)\|_{2}^{2}}{\|Ae\|_{2}}$ $\leq\|e\|_{2}\leq\frac{\beta\|e\|_{2}^{2}+\|\nabla_{\theta}d_{\Omega}(h,f)\|_{2}^{2}}{\|Ae\|_{2}},$ (27)
where $\beta=\frac{(\lambda_{\max}(A)-\lambda_{\min}(A))^{2}}{4(\lambda_{\max}(A)+\lambda_{\min}(A))}$, and $\beta$ is the upper bound of the $T(e,Ax)$.
The proof of this theorem is provided in Appendix A.3.
To rationalize the effect of parameter quantity, we assume that for scenarios where $|\theta|\gg|h|$, the gradient $\nabla_{\theta}f_{j}$ can be considered as a random sampling from the surface of a m-dimensional sphere, $B^{m}_{\epsilon}$. This assumption is grounded in several rationales:
  1. 1.
    Gradient-based iterative algorithms typically focus on reducing the magnitude of the gradient, $\|\nabla_{\theta}f_{j}\|$, without altering its direction. This leads to near-independence between the direction vectors of $\nabla_{\theta}f_{j}$ and $\nabla_{\theta}f_{i}$ for $i\neq j$.
  2. 2.
    Randomly selecting samples in each iteration of stochastic gradient descent (SGD) or batch gradient descent (Batch-GD) leads to a random gradient direction update, thereby enhancing the independence of direction vectors.
  3. 3.
    The random initialization of parameters in DNNs(Glorot and Bengio, 2010; He et al., 2015b) helps to increase the independence between parameter gradients.
  4. 4.
    Dropout(Krizhevsky et al., 2012) ensures that the loss remains relatively constant even when specific nodes in DNNs are temporarily discarded. This effectively reduces the influence of these nodes on other parameters, thereby decreasing parameter dependency.
  5. 5.
    As the number of layers between parameter $\theta_{i}$ and $\theta_{j}$ increases, the correlation between $\nabla_{\theta_{i}}f$ and $\nabla_{\theta_{j}}f$ weakens. With a sufficiently deep DNN, it is justifiable to assume that the majority of the gradients exhibit a high degree of approximate independence.
Collectively, these arguments justify the rationale behind modeling gradient sampling on the surface of $B^{m}_{\epsilon}$, when the parameter count is sufficiently large.
Based on the established Theorem15, we deduce the following corollaries with additional insights.
Corollary 16 (number of parameters)
  1. 1.
    When $|\theta|<|h|$, we have
    $\frac{\lambda_{\max}(A)\|x\|_{2}^{2}}{4}$ $\geq\|e\|_{2}\|Ae\|_{2}-\|\nabla_{\theta}d(h,f)\|_{2}^{2}$ (28)
    $\geq 0.$
  2. 2.
    When $A=aI$, it follows that:
    $\|e\|_{2}\|Ae\|_{2}-\|\nabla_{\theta}d(h,f)\|_{2}^{2}=0.$ (29)
  3. 3.
    if $|\theta|\gg|h|$, we have
    $\lim_{|\theta|\to\infty}$ $A=aI,$ (30)
    $\lim_{|\theta|\to\infty}$ $\|e\|_{2}\|Ae\|_{2}-\|\nabla_{\theta}d(h,f)\|_{2}^{2}=0,$
    where $a\in\mathbb{R}^{+}$.
The derivations for these corollaries are presented in Appendix A.4.
These corollaries emphasizes the significance of gradient independence when the number of parameters is large, as it relates to minimizing structure errors in learning. Various techniques in deep learning, including SGD, Dropout, random initialization of parameters, deeper networks, and normalizing gradient amplitudes, contribute to this independence, ultimately enhancing the learning performance. Therefore, we believe that the ability to reduce structure errors is the main reason why these techniques can improve deep learning performance.
Theorem 15 asserts that minimizing $\|\nabla_{\theta}d_{\Omega}(h,f)\|_{2}^{2}$ and $T(e,Ae)$ results in a decrease of the value of $\|e\|_{2}$, a natural progression that subsequently guides the development of the gradient structure control algorithm.
The gradient structure control algorithm is mathematically formulated as:
$\min_{\theta}\|\nabla_{\theta}d_{\Omega}(h,f)\|_{2}^{2}+\lambda\beta,$ (31)
where $\lambda$ is a non-negative hyperparameter that weights the bound of structure error $\beta$. Optimizing this objective above derives the model’s output closer to the desired $q$, guiding the convergence of parameters to the global optimum.
Gradient-based optimization methods, including SGD, Batch-GD, and Newton’s method, primarily focus on minimizing the square $L_{2}$ norm of the gradient, $\|\nabla_{\theta}d(q,f)\|_{2}^{2}$. Alternative methods such as simulated annealing and genetic algorithms can also be employed as complementary strategies to achieve this objective. As for the optimization of $\beta$, we will establish that overparameterization, flat minima, and dynamic isometry are conducive to reducing $\beta$ in the subsequent Section 8.2.
Additionally, considering the model’s expressive capacity, complex models with greater expressive power may have multiple optima where both $|\nabla_{\theta}d_{\Omega}(q,f)|_{2}^{2}=0$ and $\beta=0$ are simultaneously met. This implies that as models grow in complexity, they possess a larger number of potential global optima, which could potentially offer advantages in terms of better fit and adaptability.
7 Bounds of learning error
This section introduces the concept of uncertainty for quantifying the risk introduced by sampling and establishing bounds on the fitting error. Leveraging the triangle inequality, we derive an upper bound for the learning error in GD Learning. Given that the traditional bias-variance trade-off may fall short in fully capturing the behavior of DNNs, we propose a new principle in GD Learning, the fitting-uncertainty trade-off. This principle, derived from the bounds on learning error, offers a new perspective on the relationship between model’s generalization ability and its fitting power.
We commence by defining generalization error as the absolute discrepancy between the expected loss under the true distribution $\bar{q}$ and its estimator $q$, formalized by:
$\mathrm{gen}(\bar{q},q,\ell):=|\mathbb{E}_{Z\sim q}\ell(z)-\mathbb{E}_{Z\sim\bar{q}}\ell(z)|,$ (32)
where $\ell:\mathcal{Z}\to[0,L]$ is a loss function with a known upper bound $L$. To establish a connection between generalization error and KL divergence, we present the following lemma:
Lemma 17
$D_{KL}(q\|\bar{q})\geq\frac{2}{L^{2}}(\mathbb{E}_{q}\ell(z)-\mathbb{E}_{\bar{q}}\ell(z))^{2}$.
The proof of this result is provided in AppendixA.5.
Since $\bar{q}$ is unknown, we model it as a random variable subject to constraints on i.i.d. sample set $s^{n}$. To formally quantify the uncertainty associated with the sampled data under the estimation $q$, we introduce the quantity $C(q)=\mathbb{E}_{\bar{q}|q}D_{KL}(q\parallel\bar{q})$, where $nq\sim\mathrm{multinomial}(\bar{q})$, and $q$ is absolutely continuous with respect to $\bar{q}$, denoted by $q\ll\bar{q}$.
Subsequently, we delve into the fundamental properties of the uncertainty measure.
Proposition 18 (properties of uncertainty)
The uncertainty possesses the followings properties:
  • As the sampling size escalates, the uncertainty tends asymptotically towards zero:
    $\lim_{n\to\infty}C(q)=0$ (33)
  • If the sample data is indifferent and all existing constraints are satisfied, a more uniform $q$ implies less uncertainty.
The proof of this result is provided in AppendixA.8.
The principle of maximum entropy posits that the most probable distribution under specified constraints is that with the maximum entropy. Since the Shannon entropy, $H(q)$, is also an indicator of $q$’s uniformity, there exists a certain equivalence between the negative Shannon entropy $-H(q)$ and the uncertainty $C(q)$.
By utilizing uncertainty, we derive upper bounds for both generalization and sampling errors, as established by the subsequent theorems, reflecting the quantitative relationship between these errors and uncertainty.
Theorem 19 (Bound on generalization error)
$\Pr(\mathrm{gen}(\bar{q},q,\ell)\geq\epsilon)\leq\frac{L^{2}C(q)}{2\epsilon^{2}}.$ (34)
The proof of this result is provided in AppendixA.6.
Theorem 20 (Bound on fitting error)
If $\bar{q}$ has finite supports, then we have
$\|q-\bar{q}\|_{2}^{2}$ $\leq 2D_{KL}(q\|\bar{q})$ (35)
$\Pr(\|q-\bar{q}\|_{2}$ $\geq\epsilon)\leq\frac{2C(q)}{\epsilon^{2}}.$
The proof of this result is provided in AppendixA.7.
Let us define:
$A_{x}=\nabla_{\theta}f_{x}^{\top}\nabla_{\theta}f_{x},e_{x}=q_{Y|x}-p_{Y|x},$ (36)
$\mathrm{gen}(\bar{q}_{X},q_{X},\|p_{Y|x}-\bar{q}_{Y|x}\|_{2})=|R_{q_{X}}(p_{Y|x},$
$\bar{q}_{Y|x})-R_{\bar{q}_{X}}(p_{Y|x},\bar{q}_{Y|x})|,$
, then we have following theorem which provides a bound on the learning error.
Theorem 21 (bound on learning error)
$R_{\bar{q}_{X}}(p_{Y|x},\bar{q}_{Y|x})$ $\leq R_{q_{X}}(p_{Y|x},q_{Y|x})+R_{q_{X}}(q_{Y|x},\bar{q}_{Y|x})$ (37)
$+\mathrm{gen}(\bar{q}_{X},q_{X},\|p_{Y|x}-\bar{q}_{Y|x}\|_{2}),$
where
$R_{q_{X}}(p_{Y|x},q_{Y|x})\leq\mathbb{E}_{X}\frac{2\beta_{x}+\|\nabla_{\theta}d_{\Omega}(q_{x},f_{x})\|_{2}^{2}}{\|A_{x}e_{x}\|_{2}},$ (38)
$\Pr(\|q_{Y|x}-\bar{q}_{Y|x}\|_{2}\geq\epsilon)\leq\frac{2C(q_{Y|x})}{\epsilon^{2}},$
$\Pr(\mathrm{gen}(\bar{q}_{X},q_{X},\|p_{Y|x}-\bar{q}_{Y|x}\|)\geq\epsilon)\leq\frac{C(q_{X})}{\epsilon^{2}}.$
The proof of this theorem is provided in Appendix A.9.
In typical deterministic classification scenarios with negligible noise, the label probability distribution conditioned on the feature $x$ is given by $\bar{q}_{Y|x}=\delta_{y}$, where $y$ is the label associated with $x$. Under these conditions, the frequency distribution aligns with the true distribution, resulting in $R_{q_{X}}(\hat{q}_{Y|x},\bar{q}_{Y|x})=0$. The bound of the learning error can then be simplified to
$R_{\bar{q}_{X}}(p_{Y|x},\bar{q}_{Y|x})$ $\leq R_{q_{X}}(p_{Y|x},\hat{q}_{Y|x})+\mathrm{gen}(\bar{q}_{X},q_{X},\|p_{Y|x}-\bar{q}_{Y|x}\|,$ (39)
where the learning error is determined by the fitting error and the uncertainty $C(q_{X})$.
Within the upper bound of the learning error, the terms $\beta_{x}$, $\|\nabla_{\theta}d_{\Omega}(q_{x},f_{x})\|_{2}^{2}$, $C(q_{Y|x})$ and $C(q_{X})$ are relatively manageable. The success of GD Learning heavily relies on the four controllable terms, which are intricately linked to the choice of $q_{X}$ and $q_{Y|x}$. Therefore, the selection of appropriate $q_{X}$ and $q_{Y|x}$ is of paramount significance in GD Learning.
Theorem 18 suggests that when the sample size is significantly large relative to the entropy, fitting error is the dominant factor in learning error. Conversely, uncertainty acts as the dominant factor. Thus, we deduce the following principle for choosing the proper $q$:
Proposition 22 (fitting-uncertainty trade-off)
In scenarios where the sample size is relatively insufficient, uncertainty plays a substantial role in learning errors, favoring the use of $q$ with higher entropy over the frequency distribution $\hat{q}$. As the sample size grows to be sufficiently large, uncertainty becomes negligible, allowing us to approximate $\bar{q}$ by $\hat{q}$.
In other words, higher uncertainty corresponds to lower confidence in the data, and therefore, a greater reliance on prior knowledge. As a result, the emphasis on data fitting should be reduced. Conversely, when uncertainty is low, the data contains almost all the information on the true underlying distribution $\bar{q}$, and the model should strive for a perfect fit on the data. The bias-variance trade-off indicates that as the complexity of the model increases, bias decreases while variance increases. They cannot both disappear and rise simultaneously. However, the results of a series of experiments on DNNs appear to contradict the traditional intuition of a strict trade-off (Neal et al., 2018). Unlike the classic bias-variance trade-off used to describe the characteristics of the model itself, our proposed fitting-uncertainty trade-off describes the relationship between the model and data, as well as how they collectively determine learning errors. Our conclusion is consistent with the formal form of maximum entropy estimation which both findings suggest that, within the given constraints, objects characterized by higher entropy, or those that exhibit a smoother nature, are generally more favorable. This principle is exemplified by techniques label smoothing (Szegedy et al., 2015), Synthetic Minority Oversampling Technique (SMOTE)(Chawla et al., 2002), which fundamentally introduce a degree of smoothness into the existing data and yield positive outcomes .
8 Understanding DL from the perspective of GD Learning
8.1 Classification and regression
In this section, we delve into the classification using softmax with cross-entropy loss and regression with mean squared error (MSE) loss. By examining these two fundamental instances, we illustrate the unification of classification and regression tasks within the GD learning framework.
The combination of softmax with cross-entropy is a standard choice to train neural network classifiers. It measures the cross-entropy between the ground truth label and the output of the neural network. The expected cross-entropy loss, denoted by $R(\theta)$, is given by:
$R(\theta)$ $=-\mathbb{E}_{X\sim q_{X}}\sum_{y\in\mathcal{Y}}q_{Y|x}(y)\log\frac{e^{f_{x}(y)}}{\sum_{y^{\prime}\in\mathcal{Y}}e^{f_{x}(y^{\prime})}}$ (40)
$=-\mathbb{E}_{X\sim q_{X}}\langle q_{Y|x},f_{x}\rangle+\mathbb{E}_{X}\log{\sum_{y\in\mathcal{Y}}e^{f_{x}(y)}}$
where $f_{x}$ is the input to the softmax function. By adding an model independent item $-H(q_{Y|X})$, we can reformulate $R(\theta)$ into a standard loss function as follows:
$\mathbb{E}_{X\sim q_{X}}d_{\Omega}(q_{Y|x},f_{x})$ $=R(\theta)-H(q_{Y|X}),$ (41)
where $H(q_{Y|X})$ represents the Shannon conditional entropy, $\Omega(q_{Y|x})=D_{KL}(q_{Y|x}\|U)$, and $U$ represents the uniform distribution over $\mathcal{Y}$. Consequently, classification with softmax cross-entropy is a specific instance of GD Learning.
Regression with MSE loss is a superior and fundamental statistical analysis method, due to its high computational efficiency and mathematical interpretability. Assuming that the label $Y$ conditioned on the given feature $x$ follows multivariate normal distribution, we can express it as:
$p(y|x)=\frac{1}{(2\pi)^{k/2}|\Sigma|^{1/2}}\exp\left(-\frac{1}{2}(y-{\mu_{x}})^{\top}\Sigma^{-1}(y-{\mu_{x}})\right),$ (42)
where $k$ is the dimension of $y$, $\Sigma=cI$, $c\in\mathbb{R}^{+}$. We approximate the Dirac Delta function $\delta_{y}$ by $\lim_{a\to 0^{+}}N(y,aI)$, where $N(y,aI)$ represents the Gaussian distribution. The standard loss is then defined as $d_{\Omega}(q_{Y|x},f_{x})=D_{KL}(q_{Y|x}\|p_{Y|x})$, where $\Omega(q_{Y|x})=D_{KL}(q_{Y|x},U)$, and $U$ is the uniform distribution over $\mathcal{Y}$. This leads to:
$D_{KL}(\delta_{y}||p_{Y|x})$ $=\frac{1}{2}[\log\frac{|cI|}{|aI|}-k+(y-{\mu_{x}})^{\top}(y-{\mu_{x}})$ (43)
$+\mathrm{trance}\left\{aI\right\}],$
$\propto(y-{\mu_{x}})^{\top}(y-{\mu_{x}}).$
Consequently, the average fitting error can be expressed as:
$\mathbb{E}_{X}$ $d_{\Omega}(q_{Y|x},f_{x})=\mathbb{E}_{X}[\mathbb{E}_{Y|x}d_{\Omega}(\delta_{y},f_{x})$ (44)
$-\mathbb{E}_{Y|x}\Omega(\delta_{y})+\Omega(q_{Y|x})]$
$=\mathbb{E}_{X}[\mathbb{E}_{Y|x}D_{KL}(\delta_{y},p_{Y|x})-\mathbb{E}_{Y|x}\Omega(\delta_{y})+\Omega(q_{Y|x})]$
$\propto\mathbb{E}_{X}\mathbb{E}_{Y|x}(y-\mu_{x})^{\top}(y-\mu_{x}).$
Thus, the regression with MSE loss is equivalent to the GD learning under Gaussian prior distribution.
8.2 Deep learning from the perspective of GD Learning
In this section, we theoretically analyze the impact of overparameterization, flat minima and dynamic isometry(Saxe et al., 2013) on learning performance with in the GD Learning framework. We demonstrate that flat minima, overparameterization and dynamic isometry serve as implementation methods of gradient structure control algorithm, providing a deeper understanding of their roles in deep learning.
First, we establish the following theorem to illustrate that in classification tasks with softmax cross-entropy loss, the largest eigenvalue of Hessian matrix associated with the loss function is bounded by that of the Gram matrix $A_{x}=\nabla_{\theta}f_{x}^{\top}\nabla_{\theta}f_{x}$, highlighting the relationship between flat minima and structure error minimization.
Lemma 23
Let $A=(a_{ij})_{m\times m}$ be a symmetric matrix. Then we have
$a_{ii}\leq\lambda_{\max}(A)\leq\mathrm{trance}(A)\leq m\lambda_{\max}(A),\forall i\leq m.$ (45)
The proof of this result is provided in Appendix A.11.
Lemma 24
Let $A,B$ be non-negative definite matrices. Then we have
$\lambda_{\min}(A+B)$ $\geq\lambda_{\min}(A)+\lambda_{\min}(B)$ (46)
$\lambda_{\max}(A+B)$ $\leq\lambda_{\max}(A)+\lambda_{\max}(B).$
The proof of this result is provided in Appendix A.10.
Theorem 25
Let us define
$\mathbb{E}_{X\sim q_{X}}d_{\Omega}(q_{Y|x})=\mathbb{E}_{X\sim q_{X}}D_{KL}(q_{Y|x},p_{Y|x}),$ (47)
and assume the following equations hold after training:
$\mathbb{E}_{Y\sim p_{Y|x}}\nabla_{\theta}f_{x}(y)=0,$ (48)
$\mathbb{E}_{X\sim q_{X}}\mathbb{E}_{Y\sim q_{Y|x}(y)}\frac{\nabla_{\theta}^{2}p_{Y|x}(y)}{p_{Y|x}(y)}=0.$
Then we have:
$\lambda_{\max}(H)\leq\max_{x}\lambda_{\max}(A_{x}),$ (49)
where $H$ is the Hessian for $\mathbb{E}_{X\sim q_{X}}D_{KL}(q_{Y|x},p_{Y|x})$.
The proof of this result is provided in Appendix A.12.
In other words, when employing softmax cross-entropy loss, which is given by $R(\theta)$, minimizing $\lambda_{\max}(A_{x}),\forall x\in\mathcal{X}$, is equivalent to minimizing $\lambda_{\max}(\nabla^{2}_{\theta}R(\theta))$. When the parameter count $|\theta|$ is less than the cardinality of the label set $\mathcal{Y}$, $A_{x}$ is underdetermined, resulting in a zero minimum eigenvalue, $\lambda_{\min}(A_{x})=0$. In this scenario, referring to Corollary 16 and Theorem 21, the fitting error bound is governed by the parameter $\beta_{x}=\frac{\lambda_{\max}(A_{x})}{4}$, suggesting that minimizing $\lambda_{\max}(A_{x})$ is advantageous in reducing fitting and learning error, aligning with the flat minima theory. Therefore, when applying softmax cross-entropy loss, flat minima can be regarded as an empirical conclusion of the specific case of Theorem 21, which offers a broader perspective on the relationship between learning performance and eigenvalues.
Next, we proceed to investigate the impact of parameter count on learning performance. Theorem 16 indicates that the convergence of structure error to zero is guaranteed as the parameter count increases. In this case, the fitting error can be controlled solely by the norm of gradients. This underlies the abilities of overparameterized models to approximate global optimal solutions for non-convex optimization through gradient-based iterative algorithms, as well as the excellent performance of DNNs.
The dynamic isometry condition posits that a model exhibits favorable properties when the singular value distribution of its input-output Jacobian matrix is nearly uniform around 1. Models that satisfy this condition at initialization are often considered to possess strong characteristics, warranting high expectations for their training outcomes, as demonstrated by empirical evidence across diverse network architectures(Xiao et al., 2018; Chen et al., 2020).
It is evident that, under the dynamic isometry condition, the value of the structure error is suppressed. This is advantageous for reducing fitting errors, as demonstrated by the bound of learning error in Corollary 16. Therefore, GD Learning also provides a theoretical rationale for dynamic isometric initialization.
The preceding analysis reveals that flat minima can be viewed as an optimization technique to mitigate structure errors, while overparameterization can be interpreted as a strategy that reduces structure errors through model architecture. Dynamic isometric initialization, on the other hand, contributes to minimizing structure errors through its initialization approach. Collectively, these well-established phenomena and theories effectively address the core issue of structure error reduction. Hence, we introduce the principle of minimum structure error. The preceding analysis provides a rationale to consider this principle as fundamental to the comprehension of deep learning mechanisms. Furthermore, the definition of structure error suggests that enhancing the orthogonality among gradient vectors and ensuring the modulus values of the constraints are roughly equivalent can effectively reduce structure error. Notably, augmenting the symmetry within the neural network architecture facilitates the attainment of orthogonality among gradients and uniformity in gradient modulus values. Consequently, we speculate that enhancing the symmetry within network architectures, which encompasses strategies such as weight regulation and the normalization of input and intermediate layer features, are practical methods to enforce the minimum structure error principle. The consistency with existing theories and experimental phenomena, as well as the explanatory power of existing theories and phenomena, demonstrate the correctness and effectiveness of our proposed framework.
9 Conclusion
This study introduces a novel theoretical learning framework, GD Learning, which offers a comprehensive perspective on various machine learning and statistical tasks. GD Learning is distinguished by its emphasis on the true underlying distribution, decomposing learning errors into sampling error and fitting errors, revealing the profound influence of the sample set on learning performance. Moreover, the framework significantly incorporates prior knowledge, especially in scenarios characterized by data scarcity measured by uncertainty. We also introduce the gradient structure control algorithm for non-convex optimization objectives, which extends gradient-based methods and establishes an effective bound on learning error under its application. We demonstrate that the GD Learning unifies diverse tasks, such as classification, regression, and parameter estimation within a unified theoretical framework. By introducing the principle of minimum structure error, our framework offers a clearer insights of unanswered problems on deep learning, including the roles of flat minima, overparameterization, and dynamical isometry in learning performance, and provides a theoretical explanation for their impact. To address the limitations of the bias-variance trade-off in deep learning, we propose and prove the fitting-uncertainty trade-off theory. It highlights the importance of balancing the fitting error of existing data with the learned distribution’s uncertainty, thus offering a tailored solution for deep learning scenarios. GD Learning is theoretically capable of elucidating the theoretical rationale behind phenomena and offering a more general form. The consistency of our framework with existing theories and experimental phenomena, as well as the explanatory power of existing theories and phenomena, demonstrate the correctness and effectiveness of our proposed framework.
Acknowledgments and Disclosure of Funding
All acknowledgements go at the end of the paper before appendices and references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found on the JMLR website.
Appendix A Proofs
A.1 Proof of Theorem 13
Sine $\ell(q,p)$ is continuous, Fréchet differentiable and has a unique minimum value of zero at $p=q$, where its gradient with respect to $p$, vanishes, i.e., $\nabla_{p}\ell(q,q)=0$, we can express this gradient as
$\nabla_{p}\ell(q,p)=(p-q)^{\top}g(q,p).$ (50)
Due to the unbiasedness of minimum value, the expected loss, $\mathbb{E}_{q\sim Q}\ell(q,p)$, is strictly convex, and at its minimum $\bar{q}$, the gradient with respect to $p$ vanishes, i.e., $\nabla_{p}\mathbb{E}_{q\sim Q}\ell(q,\bar{q})=0$, which implies that $\sum_{q\in\mathcal{Q}}Q(q)(\bar{q}-q)^{\top}g(q,\bar{q})=0$, where $q\sim Q$.
Given that $Q(q)$ is an arbitrary distribution function, $g(q,p)$ must be independent with $q$, allowing us to replace $g(q,p)$ by $g(p)$.
Since $\mathbb{E}_{q\sim Q}\ell(q,p)$ is a strictly convex function, its Hessian with respect to $p$ can be derived as:
$\nabla^{2}_{p}\mathbb{E}_{q\sim Q}\ell(q,p)=\sum_{q\in\mathcal{Q}}Q(q)[g(p)+(p-q)^{\top}\nabla_{p}g(p)].$ (51)
Simplifying we get:
$\nabla^{2}_{p}\mathbb{E}_{q\sim Q}\ell(q,p)=g(p)+(p-\bar{q})\nabla_{p}g(p)>0.$ (52)
The inequality above is derived by the fact that $\nabla^{2}_{p}\mathbb{E}_{q\sim Q}\ell(q,p)$ is strictly convex.
When $p$ is set to $\bar{q}$, we have $\nabla^{2}_{p}\mathbb{E}_{q\sim Q}\ell(q,p)=g(\bar{q})>0$. Since $\bar{q}$ can take any value, it follows that for any arbitrary $p$, we get $g(p)>0$.
Let us define a strictly convex function $\Omega(p)$ with its Hessian matrix given by $\nabla^{2}_{p}\Omega(p)=g(p)$. The gradient of the loss function $\ell(q,p)$ with respect to $p$ can be expressed as follows:
$\nabla_{p}\ell(q,p)$ $=\nabla^{2}_{p}\Omega(p)(p-q)$ (53)
$=\nabla_{p}p_{\Omega}^{*}p-\nabla_{p}p_{\Omega}^{*}q$
$=\nabla_{p}(\langle p,p_{\Omega}^{*}\rangle-\Omega(p))-q^{\top}\nabla_{p}p_{\Omega}^{*}$
$=\nabla_{p}(\Omega^{*}(p_{\Omega}^{*})-\langle q,p_{\Omega}^{*}\rangle+h(q)).$
Since $\nabla_{p}\ell(q,q)=0$, we find that $h(q)=\Omega^{*}(q_{\Omega}^{*})-\langle q,q_{\Omega}^{*}\rangle=\Omega(q)$. This leads to the following expression for the loss function: $\ell(q,p)=\Omega^{*}(p_{\Omega}^{*})-\langle q,p_{\Omega}^{*}\rangle+\Omega(q)$. According the equivalence between Bregman divergence and Fenchel-Young loss 8, the loss function can also be represented as the Bregman divergence with respect to $\Omega$, denoted by $\ell(q,p)=B_{\Omega}(q,p)$.
A.2 Proof of Theorem 14
We introduce a function $F(a)=a^{2}\|x\|_{2}^{2}+\|f\|_{2}^{2}/a^{2}$ , where $a\in\mathbb{R}^{+}$. Taking the first and second derivatives of $F(a)$, we get
$\nabla_{a}F(a)$ $=\frac{d}{da}\left(a^{2}\|x\|_{2}^{2}+\frac{\|f\|_{2}^{2}}{a^{2}}\right)$ (54)
$=2a\|x\|_{2}^{2}-\frac{2\|f\|_{2}^{2}}{a^{3}},$
$\nabla^{2}_{a}F(a)$ $=\frac{d^{2}}{da^{2}}\left(a^{2}\|x\|_{2}^{2}+\frac{\|f\|_{2}^{2}}{a^{2}}\right)$
$=\frac{d}{da}\left(2a\|x\|_{2}^{2}-\frac{2\|f\|_{2}^{2}}{a^{3}}\right)$
$=2\|x\|_{2}^{2}+\frac{6\|f\|_{2}^{2}}{a^{4}}.$
The second derivative is non-negative for all $a>0$. Consequently, $F(a)$ possesses a unique globally minimum, specifically, a value of $\|x\|_{2}\|f\|_{2}$, which is attained at the critical point $a=(\|f\|_{2}^{2}/\|x\|_{2}^{2})^{1/4}$. From the relationships:
$\|ax-f/a\|_{2}^{2}$ $=\|ax\|_{2}^{2}+\|f/a\|_{2}^{2}-\langle x,f\rangle,$ (55)
$\geq\|x\|_{2}\|f\|_{2}-\langle x,f\rangle,$
$\|ax+f/a\|_{2}^{2}$ $=\|ax\|_{2}^{2}+\|f/a\|_{2}^{2}+\langle x,f\rangle,$
$\geq\|x\|_{2}\|f\|_{2}+\langle x,f\rangle,$
where the equations hold when $a=(\|f\|_{2}^{2}/\|x\|_{2}^{2})^{1/4}$, we can derive the expression for $T(x,f)$:
$T(x,f)=\|f\|_{2}\|x\|_{2}-|\langle x,f\rangle|\geq 0,$ (56)
with equality holding iif $\forall i,j,x_{i}f_{j}=x_{j}f_{i}$. Similarly, for the expected value with respected to $X$, we have
$\mathbb{E}_{X}T(e_{x},g_{x})=\mathbb{E}_{X}\|e_{x}\|_{2}\|g_{x}\|_{2}-\mathbb{E}_{X}|\langle e_{x},g_{x}\rangle|\geq 0,$ (57)
equality holds iif $\forall i,j,g_{x}(i)e_{x}(j)=g_{x}(j)e_{x}(i)$.
A.3 Proof of Theorem 15
Let us define:
$B(a)$ $:=(aI-A/a)$ (58)
$\epsilon$ $:=(\lambda_{\max}(A)-\lambda_{\min}(A))/2\lambda_{\max}(A),$
where $0\leq\epsilon\leq\frac{1}{2}$. The following relationships hold:
$\lambda_{\min}(B(a))$ $=\lambda_{\min}(A)/a-a$ (59)
$\lambda_{\max}(B(a))$ $=\lambda_{\max}(A)/a-a$
$\lambda_{\max}(B(a)^{\top}B(a))$ $=\max\{(\lambda_{\max}(A)/a-a)^{2}$
$,(a-\lambda_{\min}(A)/a)^{2}\}.$
If the condition $\lambda_{\max}(A)/a-a\geq a-\lambda_{\min}(A)/a$ holds, the maximum eigenvalue of $B(a)^{\top}B(a)$ is given by:
$\lambda_{\max}(B(a)^{\top}B(a))=(\lambda_{\max}(A)/a-a)^{2}.$ (60)
Furthermore, when
$a=\sqrt{\lambda_{\max}(A)(1-\epsilon)},$ (61)
$\lambda_{\max}(B(a)^{\top}B(a))$ attains its minimum value, which is
$\frac{\lambda_{\max}(A)\epsilon^{2}}{1-\epsilon}.$ (62)
When the inequality $\lambda_{\max}(A)/a-a<a-\lambda_{\min}(A)/a$ holds, then the maximum eigenvalue of $B(a)^{\top}B(a)$ is given by:
$\lambda_{\max}(B(a)^{\top}B(a))=(a-\lambda_{\min}(A)/a)^{2}.$ (63)
In the case $a=\sqrt{\lambda_{\max}(A)(1-\epsilon)}$, $\lambda_{\max}(B(a)^{\top}B(a))$ reaches its minimum value, which is also
$\frac{\lambda_{\max}(A)\epsilon^{2}}{1-\epsilon}.$ (64)
Since $A$ is a symmetric matrix, we have
$\lambda_{\max}(A)$ $\geq\lambda_{\min}(A)\geq 0,$ (65)
$\min_{a\in\mathbb{R}^{+}}\|ae-Ae/a\|_{2}^{2}$ $\leq\min_{a\in\mathbb{R}^{+}}\|ae+Ae/a\|_{2}^{2}.$
Consequently, we obtain that:
$T(e,Ae)$ $=\frac{1}{2}\min_{a\in\mathbb{R}^{+}}\|ae-Ae/a\|_{2}^{2}$ (66)
$=\frac{1}{2}\min_{a\in\mathbb{R}^{+}}e^{\top}B(a)^{\top}B(a)e$
$\leq\frac{1}{2}\|e\|_{2}^{2}\min_{a\in\mathbb{R}^{+}}\lambda_{\max}(B(a)^{\top}B(a))$
$=\frac{1}{2}\|e\|_{2}^{2}\frac{\lambda_{\max}(A)\epsilon^{2}}{1-\epsilon}$
$=\frac{(\lambda_{\max}(A)-\lambda_{\min}(A))^{2}\|e\|_{2}^{2}}{4(\lambda_{\max}(A)+\lambda_{\min}(A))}.$
The last equation is derived by substituting $\epsilon$ with $\frac{1}{2}(1-\frac{\lambda_{\min}(A)}{\lambda_{\max}(A)})$.
Let $\beta=\frac{(\lambda_{\max}(A)-\lambda_{\min}(A))^{2}}{4(\lambda_{\max}(A)+\lambda_{\min}(A))}$, where $\lambda_{\min}(A)\geq 0$, we have $0\leq\beta\leq\frac{\lambda_{\max}}{4}$.
Applying theorem 14 , we obtain the following inequality:
00 0 $\leq T(e,Ae)$ (67)
$=\|e\|_{2}\|Ae\|_{2}-\|\nabla_{\theta}d_{\Omega}(q,f)\|_{2}^{2}$
$\leq\beta\|e\|_{2}^{2}.$
Consequently, we have
$\frac{\|\nabla_{\theta}d_{\Omega}(q,f)\|_{2}^{2}}{\|Ae\|_{2}}\leq\|e\|_{2}\leq\frac{\beta\|e\|_{2}^{2}+\|\nabla_{\theta}d_{\Omega}(q,f)\|_{2}^{2}}{\|Ae\|_{2}}$ (68)
A.4 Proof of Corollary 16
When $|\theta|<|h|$, the matrix $A$ is underdetermined, resulting in $\lambda_{\min}(A)=0$. Leveraging the theorem 14, we obtain:
$\lambda_{\max}(A)\|x\|_{2}^{2}/4\geq T(e,Ae)=\|e\|_{2}\|Ae\|_{2}-\|\nabla_{\theta}F(f)\|_{2}^{2}\geq 0.$ (69)
In the special case where $A=aI$, the inequality above simplifies to $T(e,Ae)=0$, thus $\|e\|_{2}\|Ae\|_{2}-\|\nabla_{\theta}F(f)\|_{2}^{2}=0$.
[…]
$\lim_{|\theta|\to\infty}\|e\|_{2}={\|\nabla_{\theta}F(f)\|_{2}^{2}}/\|Ae\|_{2}.$ (70)
A.5 Proof of Lemma 17
Firstly, let’s derive the upper bound of a as follows:
$|\mathbb{E}_{q}$ $\ell-\mathbb{E}_{\bar{q}}\ell|=|\sum_{z\in\mathcal{Z}}(q(z)-\bar{q}(z))\ell(z)|$ (71)
$=|\sum_{z\in\mathcal{Z}}(\bar{q}(x)-q(x))(\ell(z)-\frac{L}{2})+\frac{L}{2}(\bar{q}(z)-q(z))|$
$\leq\sum_{z\in\mathcal{Z}}|\bar{q}(z)-q(z)||\ell(z)-\frac{L}{2}|$
$\leq\frac{L}{2}\|q-\bar{q}\|_{1}$
Employing Pinsker’s inequality (Csiszar, 1967; Kullback, 1967), we have:
$D_{KL}(q\|\bar{q})\geq\frac{1}{2}\|q-\bar{q}\|_{1}^{2}.$ (72)
Combining equation (71) and equation (72) gives us:
$D_{KL}(q\|\bar{q})\geq\frac{2}{L^{2}}|\mathbb{E}_{q}\ell(z)-\mathbb{E}_{\bar{q}}\ell(z)|^{2}$ (73)
A.6 Proof of Theorem 19
According to Lemma 17, we have
$D_{KL}(q\|\bar{q})\geq\frac{2}{L^{2}}\mathrm{gen}(\bar{q},q,\ell)^{2},$ (74)
where $L=\max_{x}\ell(x)$. Consequently, we obtain:
$\Pr(D_{KL}(q\|\bar{q})\geq t)$ $\geq\Pr(\frac{2}{L^{2}}\mathrm{gen}(\bar{q},q,\ell)^{2}\geq t)$ (75)
$=\Pr(\mathrm{gen}(\bar{q},q,\ell)^{2}\geq L^{2}t/2)$
$=\Pr(\mathrm{gen}(\bar{q},q,\ell)\geq L\sqrt{t/2}).$
Let $\epsilon$ be defined as $\epsilon=L\sqrt{t/2}$. By applying lemma 11, we obtain the subsequent inequality:
$\Pr(\mathrm{gen}(\bar{q},f)\geq\epsilon)$ $\leq\Pr(D_{KL}(q\|\bar{q})\geq\frac{2\epsilon^{2}}{L^{2}})$ (76)
$\leq\frac{L^{2}\mathbb{E}_{\bar{q}|q}D_{KL}(q\|\bar{q})}{2\epsilon^{2}}.$
A.7 Proof of Theorem 20
Let $e$ be a vector of dimension $k$, then we have
$\|e\|_{1}^{2}$ $=\sum_{i=1}^{k}\sum_{j=1}^{k}\|e(i)\|e(j)\|$ (77)
$\geq\sum_{i=1}^{k}\|e(i)\|^{2}$
$=\|e\|_{2}^{2}.$
Employing Pinsker’s inequality (72), we have:
$D_{KL}(q\|\bar{q})\geq\frac{1}{2}\|\bar{q}-q\|_{1}^{2}.$ (78)
Since $\bar{q}$ has finite supports, combining the inequality (77) and inequality (78), we have
$D_{KL}(q\|\bar{q})\geq\frac{1}{2}\|q-\bar{q}\|_{2}^{2}.$ (79)
Applying lemma 11, we obtain the following result:
$\Pr(\|q-\bar{q}\|_{2}\geq\epsilon)$ $\leq\Pr(2D_{KL}(q\|\bar{q})\geq\epsilon^{2})$ (80)
$\leq\frac{2\mathbb{E}_{\bar{q}|q}D_{KL}(q\|\bar{q})}{\epsilon^{2}}.$
A.8 Proof of Proposition 18
According to the Bayesian theorem and Sanov’s theorem (Cover, 1999), we have
[…]
Combining the above equations, as $n$ approaches infinity, we obtain:
$\lim_{n\to\infty}\Pr(\bar{q}=q|q)$ $=1$ (82)
$\lim_{n\to\infty}\mathbb{E}_{\bar{q}|q}D_{KL}(q\|\bar{q})$ $=0.$
Now, let’s find the lower bound of $C(q)$:
$C(q)$ $=-H(q)-\mathbb{E}_{\bar{q}|q}\mathbb{E}_{Z\sim q}\log\bar{q}(z)$ (83)
$=-H(q)-\mathbb{E}_{Z\sim q}\mathbb{E}_{\bar{q}|q}\log\bar{q}(z)$
$\geq-H(q)-\mathbb{E}_{Z\sim q}\log\mathbb{E}_{\bar{q}|q}\bar{q}(z),$
where the final inequality is derived from the Jensen’s inequality, as $-\log(\cdot)$ is a convex function. Given that $\sum_{Z}\mathbb{E}_{\bar{q}|q}\bar{q}(z)=1$, we have
$C(q)\geq D_{KL}(q\|\mathbb{E}_{\bar{q}|q}\bar{q}).$ (84)
Then the upper bound of $C(q)$ is derived as follows:
$C(q)$ $=-H(q)-\mathbb{E}_{\bar{q}|q}\mathbb{E}_{Z\sim q}\log\bar{q}(z)$ (85)
$=-H(q)+\mathbb{E}_{\bar{q}|q}\mathbb{E}_{Z\sim q}[1+\frac{1-\bar{q}(z)}{\bar{q}(z)}]$
$\leq-H(q)+\mathbb{E}_{\bar{q}|q}\mathbb{E}_{Z\sim q}\frac{1-\bar{q}(z)}{\bar{q}(z)},$
where the final inequality is base on $\log(1+x)\leq x$. Combine the formula (84) and (85), we have
$D_{KL}(q\|\mathbb{E}_{\bar{q}|q}\bar{q})$ $\leq C(q)$ (86)
$\leq-H(q)+\mathbb{E}_{\bar{q}|q}\mathbb{E}_{Z\sim q}\frac{1-\bar{q}(z)}{\bar{q}(z)}$
$\leq H(q)+\mathbb{E}_{\bar{q}|q}\sum_{Z}\frac{q_{\max}}{\bar{q}(z)}-1.$
Because $\bar{q}$ is an arbitrary distribution on $\mathcal{Z}$ , when $n$ is insufficient large, the conditional probability $\Pr(\bar{q}=q|q)$ fails to approximate to 1, causing $\mathbb{E}_{\bar{q}|q}\bar{q}$ to be a relatively uniform. Consequently, a more uniform $q$ often implies a smaller lower $D_{KL}(q\|\mathbb{E}_{\bar{q}|q}\bar{q})$, i.e., smaller bound of $C(q)$. The fact that $q_{\max}$ influences the upper bound also highlights the advantage of a uniform $q$ in minimizing the $C(q)$. Thus, under the condition of satisfying existing constraints, a more uniform $q$ implies lower uncertainty.
A.9 Proof of Theorem 21
Employing the triangle inequality of norms, we obtain the following relationship:
$R_{q_{X}}(p_{Y|x},\bar{q}_{Y|x})$ $\leq R_{q_{X}}(p_{Y|x},q_{Y|x})+R_{q_{X}}(q_{Y|x},\bar{q}_{Y|x}).$ (87)
Based on the definition of $\mathrm{gen}(\bar{q}_{X},q_{X},\|p_{Y|x}-\bar{q}_{Y|x}\|_{2})$, we have
$R_{\bar{q}_{X}}(p_{Y|x},\bar{q}_{Y|x})$ $\leq R_{q_{X}}(p_{Y|x},\bar{q}_{Y|x})$ (88)
$+\mathrm{gen}(\bar{q}_{X},q_{X},\|p_{Y|x}-\bar{q}_{Y|x}\|_{2}).$
Combining the above inequality and the inequality (87), we obtain
$R_{\bar{q}_{X}}(p_{Y|x},\bar{q}_{Y|x})$ $\leq R_{q_{X}}(p_{Y|x},q_{Y|x})+R_{q_{X}}(q_{Y|x},\bar{q}_{Y|x})$ (89)
$+\mathrm{gen}(\bar{q}_{X},q_{X},\|p_{Y|x}-\bar{q}_{Y|x}\|_{2}).$
Given that $\|p_{Y|x}-\bar{q}_{Y|x}\|_{2}\leq\sqrt{2}$, by Theorem19, we have
$\Pr(\mathrm{gen}(\bar{q}_{X},q_{X},\|p_{Y|x}-\bar{q}_{Y|x}\|_{2}\geq\epsilon)\leq\frac{C(q_{X})}{\epsilon^{2}}.$ (90)
According to Theorem 20, we have:
$\Pr(\|q_{Y|x}-\bar{q}_{Y|x}\|_{2}\geq\epsilon)\leq\frac{2C(q_{Y|x})}{\epsilon^{2}}.$ (91)
Invoking Theorem 15, we have:
$\frac{\|\nabla_{\theta}d_{\Omega}(q_{Y|x},f_{x})\|_{2}^{2}}{\|A_{x}e_{x}\|_{2}}\leq\|e_{x}\|_{2}\leq\frac{2\beta_{x}+\|\nabla_{\theta}d_{\Omega}(q_{Y|x},f_{x})\|_{2}^{2}}{\|A_{x}e_{x}\|_{2}},$ (92)
where $\beta_{x}=\frac{(\lambda_{\max}(A_{x})-\lambda_{\min}(A_{x}))^{2}}{4(\lambda_{\max}(A_{x})+\lambda_{\min}(A_{x}))}$.
Taking the expectation of above inequality respect to $X$, we derive:
$\mathbb{E}_{X}\frac{\|\nabla_{\theta}d_{\Omega}(q_{x},f_{x})\|_{2}^{2}}{\|A_{x}e_{x}\|_{2}}$ $\leq R_{q_{X}}(p_{Y|x},q_{Y|x})$ (93)
$\leq\mathbb{E}_{X}\frac{2\beta_{x}+\|\nabla_{\theta}d_{\Omega}(q_{x},f_{x})\|_{2}^{2}}{\|A_{x}e_{x}\|_{2}}.$
Combining the inequalities (93),(90),(80), we obtain
$R_{\bar{q}_{X}}(p_{Y|x},\bar{q}_{Y|x})\leq R_{q_{X}}(p_{Y|x},q_{Y|x})+R_{q_{X}}(q_{Y|x},\bar{q}_{Y|x})$ (94)
$+\mathrm{gen}(\bar{q}_{X},q_{X},\|p_{Y|x}-\bar{q}_{Y|x}\|_{2}),$
$\mathbb{E}_{X}\frac{\|\nabla_{\theta}d_{\Omega}(q_{x},f_{x})\|_{2}^{2}}{\|A_{x}e_{x}\|_{2}}\leq R_{q_{X}}(p_{Y|x},q_{Y|x})$
$\leq\mathbb{E}_{X}\frac{2\beta_{x}+\|\nabla_{\theta}d_{\Omega}(q_{x},f_{x})\|_{2}^{2}}{\|A_{x}e_{x}\|_{2}},$
$\Pr(\mathrm{gen}(\bar{q}_{X},q_{X},\|p_{Y|x}-\bar{q}_{Y|x}\|_{2}\geq\epsilon)\leq\frac{C(q_{X})}{\epsilon^{2}},$
$\Pr(\|q_{Y|x}-\bar{q}_{Y|x}\|_{2}\geq\epsilon)\leq\frac{2C(q_{Y|x})}{\epsilon^{2}}.$
A.10 Proof of Lemma 24
Since $x^{\top}Ax\geq\lambda_{\min}(A)x^{\top}x$, it follows that
$x^{\top}(A-\lambda_{\min}(A)I)x\geq 0,$ (95)
and $A-\lambda_{\min}(A)I$ is non-negative definite matrix. Therefore, we have
$A-\lambda_{\min}(A)I\succeq 0$ (96)
$B-\lambda_{\min}(B)I\succeq 0$
$A-\lambda_{\min}(A)I+B-\lambda_{\min}(B)I\succeq 0.$
Let $x^{\prime}$ satisfies the equality $(A+B)x^{\prime}=\lambda_{\min}(A+B)x^{\prime}$, then we have
$(A-$ $\lambda_{\min}(A)I+B-\lambda_{\min}(B)I)x^{\prime}$ (97)
$=(\lambda_{\min}(A+B)-\lambda_{\min}(A)-\lambda_{\min}(B))x^{\prime}.$
Therefore $\lambda_{\min}(A+B)-\lambda_{\min}(A)-\lambda_{\min}(B)$ is the eigenvalue of $A-\lambda_{\min}(A)I+B-\lambda_{\min}(B)I$. As $A-\lambda_{\min}(A)I+B-\lambda_{\min}(B)I$ is non-negative definite matrix, thus we have $\lambda_{\min}(A+B)-\lambda_{\min}(A)-\lambda_{\min}(B)\geq 0$.
Because $x^{\top}A\leq\lambda_{\max}(A)x^{\top}x$, then $x^{\top}(A-\lambda_{\max}(A)I)x\leq 0$, thus $\lambda_{\max}(A)I-A$ is non-negative definite matrix. Analogously, we can obtain $\lambda_{\max}(A+B)-\lambda_{\max}(A)-\lambda_{\max}(B)\leq 0$.
A.11 Proof of Lemma 23
Given a symmetric matrix $A=(a_{ij})_{m\times m}$, we know that $\lambda_{\min}(A)\geq 0$. Since the sum of eigenvalue is equal to $\mathrm{trance}(A)$, so we have
$\lambda_{\max}(A)\leq\mathrm{trance}(A)\leq m\lambda_{\max}(A).$ (98)
By the Courant-Fisher min-max theorem, the maximum eigenvalue is given by:
$\lambda_{\max}(A)=\max_{x\in\mathbb{R}^{m},\|x\|=1}x^{\top}Ax.$ (99)
Let $e_{i}$ be a standard basis (“one-hot”) vector, then we have
$a_{ii}=e_{i}^{\top}Ae_{i}\leq\lambda_{\max}(A)$ (100)
A.12 Proof of Theorem 25
Given the relationship: $\mathbb{E}_{X\sim q_{X}}d_{\Omega}(q_{Y|x},f_{x})=\mathbb{E}_{X\sim q_{X}}D_{KL}(q_{Y|x}\|p_{Y|x})$, we can deduce $p_{Y|x}(y)=\frac{e^{f_{x}(y)}}{Z_{x}}$, $Z_{x}$ is the partition function, defined as $Z_{x}=\sum_{y\in\mathcal{Y}}e^{f_{x}(y)}$. Consequently, the gradient of the log-probability with respect to the model parameter $\theta$ decomposes as follows:
$\nabla_{\theta}\log p_{Y|x}(y)=\nabla_{\theta}f_{x}(y)-\nabla_{\theta}\log Z_{x},$ (101)
The second gradient term, $\nabla_{\theta}\log Z_{x}$ can be rewritten as an expected value, as follows:
$\nabla_{\theta}\log Z_{x}$ $=Z_{x}^{-1}\nabla_{\theta}\sum_{y\in\mathcal{Y}}e^{f_{x}(y)}$ (102)
$=Z_{x}^{-1}\sum_{y\in\mathcal{Y}}e^{f_{x}(y)}\nabla_{\theta}f_{x}(y)$
$=\mathbb{E}_{Y\sim p_{Y|x}}\nabla_{\theta}f_{x}(y).$
According to the above equality, we have
$\mathbb{E}_{Y\sim q_{Y|x}}\left\{\nabla_{\theta}\log p_{Y|x}(y)\right\}\left\{\nabla_{\theta}\log p_{Y|x}(y)\right\}^{\top}$ (103)
$=\mathbb{E}_{Y\sim q_{Y|x}}\{\nabla_{\theta}f_{x}(y)-\nabla_{\theta}\log Z_{x}\}\{\nabla_{\theta}f_{x}(y)-\nabla_{\theta}\log Z_{x}\}^{\top}$
$=\mathbb{E}_{Y\sim q_{Y|x}}\{\nabla_{\theta}f_{x}(y)\nabla_{\theta}f_{x}(y)^{\top}\}$
$-2\nabla_{\theta}\log Z_{x}\mathbb{E}_{Y\sim q_{Y|x}}\nabla_{\theta}f_{x}(y)+\nabla_{\theta}\log Z_{x}\nabla_{\theta}\log Z_{x}^{\top}$
$=\mathbb{E}_{Y\sim q_{Y|x}}\{\nabla_{\theta}f_{x}(y)\nabla_{\theta}f_{x}(y)^{\top}\}+F_{x}$
where
$F_{x}=-2\nabla_{\theta}\log Z_{x}\mathbb{E}_{Y\sim q_{Y|x}}\nabla_{\theta}f_{x}(y)+\nabla_{\theta}\log Z_{x}\nabla_{\theta}\log Z_{x}^{\top}.$ (104)
The Hessian for $D_{KL}(q_{Y|x},p_{Y|x})$ is equal to compute with the second derivative of negative log-likelihood $-\log p_{Y|x}(y)$, which can be expressed as:
$\nabla^{2}_{\theta}D_{KL}(q_{Y|x},p_{Y|x})=\mathbb{E}_{Y\sim q_{Y|x}}[-\nabla^{2}_{\theta}\log p_{Y|x}(y)].$ (105)
Next, we expand the second derivative of it as follows:
-- $\nabla_{\theta}^{2}\log p_{Y|x}(y)=-\nabla_{\theta}\left[\frac{\nabla_{\theta}p_{Y|x}(y)}{p_{Y|x}(y)}\right]$ (106)
$=-\frac{p_{Y|x}(y)\nabla_{\theta}^{2}p_{Y|x}(y)-\left\{\nabla_{\theta}p_{Y|x}(y)\right\}\left\{\nabla_{\theta}p_{Y|x}(y)\right\}^{\top}}{p_{Y|x}(y)^{2}}$
$=-\frac{\nabla_{\theta}^{2}p_{Y|x}(y)}{p_{Y|x}(y)}+\left\{\frac{\nabla_{\theta}p_{Y|x}(y)}{p_{Y|x}(y)}\right\}\left\{\frac{\nabla_{\theta}p_{Y|x}(y)}{p_{Y|x}(y)}\right\}^{\top}$
$=-\frac{\nabla_{\theta}^{2}p_{Y|x}(y)}{p_{Y|x}(y)}+\left\{\nabla_{\theta}\log p_{Y|x}(y)\right\}\left\{\nabla_{\theta}\log p_{Y|x}(y)\right\}^{\top}.$
Then, we apply this expanded second derivative to (105) which can be restated as follows:
$\nabla^{2}_{\theta}D_{KL}(q_{Y|x},p_{Y|x})=\mathbb{E}_{Y\sim q_{Y|x}(y)}\left[-\frac{\nabla_{\theta}^{2}p_{Y|x}(y)}{p_{Y|x}(y)}\right]$ (107)
$+\mathbb{E}_{Y\sim q_{Y|x}(y)}\left[\left\{\nabla_{\theta}\log p_{Y|x}(y)\right\}\left\{\nabla_{\theta}\log p_{Y|x}(y)\right\}^{\top}\right]$
$=C_{x}+B_{x}+F_{x}.$
where
$B_{x}$ $=\mathbb{E}_{Y\sim p_{Y|x}}\nabla_{\theta}f_{x}(y)\nabla_{\theta}f_{x}(y)^{\top},$ (108)
$C_{x}$ $=\mathbb{E}_{Y\sim q_{Y|x}(y)}\left[-\frac{\nabla_{\theta}^{2}p_{Y|x}(y)}{p_{Y|x}(y)}\right].$
Since $\nabla_{\theta}f_{x}(y)\nabla_{\theta}f_{x}(y)^{\top}$ is matrix of rank 1, its eigenvalues are
$\lambda_{\max}(\nabla_{\theta}f_{x}(y)\nabla_{\theta}f_{x}(y)^{\top})$ $=\nabla_{\theta}f_{x}(y)^{\top}\nabla_{\theta}f_{x}(y)$ (109)
$\lambda_{\min}(\nabla_{\theta}f_{x}(y)\nabla_{\theta}f_{x}(y)^{\top})$ $=0$
According to Lemma 24, we can bounded the extreme eigenvalues of $B_{x}$ as follows:
$\lambda_{\max}(B_{x})$ $\leq\mathbb{E}_{Y\sim p_{Y|x}}\lambda_{\max}(\nabla_{\theta}f_{x}(y)\nabla_{\theta}f_{x}(y)^{\top})$ (110)
$\leq\max_{y}\lambda_{\max}(\nabla_{\theta}f_{x}(y)\nabla_{\theta}f_{x}(y)^{\top})$
$=\max_{y}\nabla_{\theta}f_{x}(y)^{\top}\nabla_{\theta}f_{x}(y)$
$\lambda_{\min}(B_{x})$ $\geq\mathbb{E}_{Y\sim p_{Y|x}}\lambda_{\min}(\nabla_{\theta}f_{x}(y)\nabla_{\theta}f_{x}(y)^{\top})$
$\geq\min_{y}\lambda_{\min}(\nabla_{\theta}f_{x}(y)\nabla_{\theta}f_{x}(y)^{\top})$
$=0.$
Because $A_{x}=\nabla_{\theta}f_{x}^{\top}\nabla_{\theta}f_{x}$ is a symmetric matrix, according to the inequality (110) and lemma23 , we have:
$\lambda_{\max}(B_{x})\leq\max_{y}\nabla_{\theta}f_{x}(y)^{\top}\nabla_{\theta}f_{x}(y)$ $\leq\lambda_{\max}(A_{x}).$ (111)
The Hessian for $\mathbb{E}_{X\sim q_{X}}D_{KL}(q_{Y|x},p_{Y|x})$, can be expressed as:
$H=\mathbb{E}_{X\sim q_{X}}B_{x}+\mathbb{E}_{X\sim q_{X}}C_{x}+\mathbb{E}_{X\sim q_{X}}F_{x}.$ (112)
Since
$\mathbb{E}_{Y\sim p_{Y|x}}\nabla_{\theta}f_{x}(y)=0,$ (113)
$\mathbb{E}_{X\sim q_{X}}\mathbb{E}_{Y\sim q_{Y|x}(y)}\frac{\nabla_{\theta}^{2}p_{Y|x}(y)}{p_{Y|x}(y)}=0,$
, we obtain
$H$ $=\mathbb{E}_{X\sim q_{X}}B_{x},$ (114)
$\lambda_{\max}(H)$ $\leq\max_{x}\lambda_{\max}(B_{x})\leq\max_{x}\lambda_{\max}(A_{x})$
Remainder omitted in this sample. See http://www.jmlr.org/papers/ for full paper.